### v1

# HDC: Hierarchical Semantic Decoding with Counting Assistance for Generalized Referring Expression Segmentation

Abstract
The newly proposed Generalized Referring Expression Segmentation (GRES) amplifies the formulation of classic RES by involving multiple/non-target scenarios. Recent approaches focus on optimizing the last modality-fused feature which is directly utilized for segmentation and object-existence identification. However, the attempt to integrate all-grained information into a single joint representation is impractical in GRES due to the increased complexity of the spatial relationships among instances and deceptive text descriptions. Furthermore, the subsequent binary target justification across all referent scenarios fails to specify their inherent differences, leading to ambiguity in object understanding. To address the weakness, we propose a Hierarchical Semantic Decoding with Counting Assistance framework (HDC). It hierarchically transfers complementary modality information across granularities, and then aggregates each well-aligned semantic correspondence for multi-level decoding. Moreover, with complete semantic context modeling, we endow HDC with explicit counting capability to facilitate comprehensive object perception in multiple/single/non-target settings. Experimental results on gRefCOCO, Ref-ZOM, R-RefCOCO, and RefCOCO benchmarks demonstrate the effectiveness and rationality of HDC which outperforms the state-of-the-art GRES methods by a remarkable margin. Code will be available here.
1 Introduction
Referring Expression Segmentation (RES) is one of the fundamental tasks in visual-linguistic understanding, which aims at localizing the referred object in an image with a natural language description. Despite achieving great progress in recent years [10, 33, 6, 21, 24], the simplification of RES formulation, where one object must be included in a sentence, still constrains its potential in various scenarios. Consequently, such limitation encourages the emergence of Generalized Referring Expression Segmentation (GRES). Different from the one-to-one setting in classic RES, GRES supports extra multiple-target and empty-target cases, rendering it more suitable in real-world applications.
However, it poses new challenges with the additional generalized settings. The multiple-target referent necessitates more precise semantic clues to handle complicated spatial-text relationships between instances for arbitrary location segmentation. On the other hand, the empty-target one highlights the significance of contextual information in mitigating undesirable responses from similar objects. Therefore, an intuitive approach is to endow prevalent RES paradigms [33, 6, 7] with new capability. For instance, ReLA [20] introduces a large-scale generalized dataset gRefCOCO and utilizes spatial-aware queries to transform GRES into a related region retrieval task. Meanwhile, DMMI [11] and RefSegformer [30] enhance modality interaction in the early-fusion process and incorporate new benchmarks Ref-ZOM and R-RefCOCO respectively, to realize robust segmentation.
Figure 1: Illustration of different decoding processes. Previous GRES methods attempt to integrate all-grained context into one joint feature which is directly for mask and object-existence prediction (a), exhibiting vulnerability to challenges of multiple/non-target scenarios (c). Instead of relying on the single representation, we perform hierarchical semantic decoding with multi-level visual-linguistic correspondence aggregation for robust segmentation. Moreover, in contrast to simple binary classification, we empower HDC with counting ability to promote overall object awareness (b).
Although these methods have made advancements in GRES, they still have intrinsic weaknesses in two aspects: 1) They are limited by the bottleneck of capturing complete associations between visual and linguistic contents. As Fig. 1 shows, they strive to optimize the last modality-fused feature which is directly for final mask prediction. However, the extra complexity of text descriptions and geometric relationships among instances in GRES impedes the integration of all-grained contexts into one joint representation. Without complete modality alignment at each granularity, it potentially leads to incorrect or incomplete object response. 2) A simple binary classification head for object-existence identification not only occurs vulnerability to deceptive instances with accumulated bias in single representation but also incurs ambiguity for ignoring inherent differences of one-one/many referents.
Therefore, the significant problem lies in how to excavate the potential of multi-grained semantic correspondence between visual and linguistic when decoding, as well as fully facilitate the comprehensive object perception by considering the specificity of each scenario. Motivated by that, we present a Hierarchical Semantic Decoding with Counting Assistance framework (HDC). In HDC, a Hierarchical Semantic Decoding module (HSD) progressively transfers complementary information between modalities across different granularities to formulate intricate instance-linguist correlation hierarchies. Instead of leveraging single joint features, HSD utilizes multi-level semantic aggregation for robust segmentation. With hierarchical semantic-context alignment, we further introduce Multi-Adaptive Object Counting module (MAC) on HSD to empower the model with explicit counting ability, which improves overall object awareness under each referring setting.
We conduct extensive experiments on three popular GRES benchmarks, i.e., gRefCOCO [20], Ref-ZOM [11] and R-RefCOCO [30] and RES datasets RefCOCO/+/g [15, 27] to validate the effectiveness of our methods. Results show that HDC sufficiently meets the challenges of GRES and addresses the limitations of existing methods, which notably outperforms them for all benchmarks.
Overall, our contributions are summarized as follows:
  • We present a Hierarchical Semantic Decoding with Counting Assistance framework (HDC) for GRES to boost the multi-grained comprehension of complex referring semantic associations at different levels. In HDC, a Hierarchical Semantic Decoding module (HSD) is designed to progressively transfer mutual modality complementary information across granularities, and utilize each well-aligned semantic correspondence for multi-level decoding.
  • Built on HSD, we further introduce Multi-Adaptive Object Counting module (MAC) to facilitate comprehensive object perception by assigning explicit counting capability, which seamlessly adapts to the specificity of multiple/single/non-target scenarios. To our best knowledge, we are the first to bring object counting into GRES.
  • Without bells and whistles, our approach outperforms all existing state-of-the-art methods by a remarkable margin under fair comparison, e.g., $+4.7\%$ gIoU in gRefCOCO, $+6.7\%$ rIoU in R-RefCOCO and $+5.3\%$ Acc. in Ref-ZOM.
2 Related Work
Referring Expression Segmentation.
Referring Expression Segmentation (RES) aims to localize a single object referred by a natural language description, which is first introduced by Hu et al. [10]. Conventional approaches [10, 22, 1, 12, 7] typically utilize convolution-based operations as the cross-modal fusion to generate segmentation masks. To overcome the weakness of insufficiency in modeling visual-text relation that prior works suffer, recent studies [33, 6, 21, 24, 16] design advanced attention-based mechanism [29] to facilitate multi-modal interaction.
Since previous works are restricted by pre-defined rules that one sentence must match an object, ReLA [20] proposes the Generalized Referring Expression Segmentation (GRES) task, which additionally involves both empty-target and multiple-target scenarios. Moreover, DMMI [11] introduces a new benchmark and baseline to realize beyond one target segmentation. Similarly, RefSegformer [30] endows the transformer-based model with empty-target sentence discrimination, achieving robust segmentation. However, they seek to integrate all-grained contents into a single joint representation for mask prediction, which is impractical in GRES. In contrast, we aggregate visual-linguistic semantic clues in hierarchical structures for more generalized segmentation.
Object Counting.
Object counting is to predict the number of conditional objects within an image, which can be divided into detection-based and regression-based approaches. The detection-based methods [13, 8, 9] obtain the number of instances by proposals localization and ranking while the regression-based ones [3, 18, 28] generates predicted counts via density map integration. Different from the classic object counting mechanism, our proposed Multi-Adaptive Object Counting module performs at the query level. By considering various scenarios, i.e., multiple/single/non-target referents, our model achieves comprehensive object perception under each situation.
3 Method
3.1 Overview
Figure 2: Overview of HDC. The model takes an image with the referring expression as input. After the encoding process, the Hierarchical Semantic Decoding module (HSD) transfers mutual modality complementary information hierarchically for multi-grained alignment and then aggregates each to perform multi-level decoding for final segmentation. Moreover, with well-aligned semantic contexts, we introduce Multi-Adaptive Object Counting module (MAC) to fully promote comprehensive object perception by the category-specific counting ability in multiple/single/non-target scenarios.
Fig. 2 depicts the overall pipeline of our HDC. Given an image $\mathcal{I}$, the visual backbone, e.g., Swin-Base [25] is to extract a set of vision features $\mathcal{V}_{i}\in\mathrm{R}^{D^{V}\times H_{i}\times W_{i}}$, $i\in\{1,2,3,4\}$, where $H_{i}$ and $W_{i}$ denote the width and height of each scale of feature maps respectively. $D^{V}$ signifies the channel dimension. $\mathcal{V}_{i}$ is further encoded as mask features $\mathcal{V}^{m}_{i}$ by deformable attention [35]. Simultaneously, the textual expression $\mathcal{E}$ with $T$ words is sent into a transformer-based [5] language encoder which generates the word-level embedding $\mathcal{L}^{w}\in\mathbb{R}^{T\times D^{L}}$ and sentence-level embedding $\mathcal{L}^{s}\in\mathbb{R}^{D^{L}}$. Considering that $\mathcal{L}^{w}$ contains fine-grained descriptions while $\mathcal{L}^{s}$ expresses the general information, language query $\mathcal{L}^{O}\in\mathbb{R}^{N\times D^{L}}$ is obtained via the concatenation of them. Meantime, $\mathcal{L}^{O}$ and $\mathcal{V}^{m}_{i}$ are sent into the Hierarchical Semantic Decoding (HSD) module, which generates corresponding object segmentation $\mathcal{M}\in\mathbb{R}^{H\times W}$ by the progressively mutual modality complementary across multi-granularity and multi-level semantic aggregation. In addition, Multi-Adaptive Object Counting (MAC) is designed to promote the precise object perception of different scenarios by category-specific counting ability. We elaborate on each module in the following sections.
3.2 Hierarchical Semantic Decoding Module
The extended formulation of multiple-target and empty-target referents in GRES brings significant challenges since the former necessities a more precise cross-modal understanding of the complex association between objects, e.g., different domains, geometrical relationships while the latter highlights long-range contextual information for mitigating interference from similar deceptive objects. Previous GRES methods [20, 11, 30] struggle to optimize the last modality-fused features for the subsequent mask prediction, ignoring the significance of explicit modality alignment at each granularity. While it is acceptable in classic RES, the extremely intricate referring semantics between instances and text in GRES hinder the integration of all-grained modality correspondence in one representation, thus incurring misinterpretation. To address the issues, we propose a Hierarchical Semantic Decoding module to perform mutual modality calibration at each granularity and multi-level semantic integration for robust segmentation. Specifically, we utilize cascaded Semantic To Mask/Query modules to progressively transfer complementary information between modalities, which captures the high-level visual-linguistic correlation at each hierarchy. Then the subsequent Multi-level Mask Aggregation is to perform semantic aggregation for object-aware segmentation.
Semantic To Mask Module.
The core idea of STM is to perform bi-directional modality complementary to generate the fine-grained semantic map at each scale of visual features, which is further utilized in the multi-level decoding process. As Fig. 3 shows, the visual feature $\mathcal{V}^{m}_{i}$ and the language query $\mathcal{L}^{O}$ are concurrently input in STM. To be specific, taking $\mathcal{V}^{m}_{1}$ as an example, we firstly leverage fully connected layers to map visual features and language query into a shared joint space $D$, respectively. The mapped representations are further denoted as $\mathcal{V}$ and $\mathcal{L}$ for clarity. Then, we derive the coarse semantic map $\mathcal{A}$ by attention mechanism and it is applied in visual-linguistic interaction to align information flow between modalities. The preliminary modality complementary highlights the corresponding object activation of language-guided visual features, while simultaneously relieving the potential noise in visual-oriented textual representations. The above process can be formulated as:
$\mathcal{A}=\frac{1}{\sqrt{D}}\left(\mathcal{L}\mathcal{W}^{k}_{L}\right)\left(\mathcal{V}\mathcal{W}^{k}_{V}\right)^{T},$ (1)
$\mathcal{F}^{L\sim V}=\mathrm{softmax}(\mathcal{A})\times(\mathcal{V}\mathcal{W}^{v}_{V}),$ (2)
$\mathcal{F}^{V\sim L}=\mathrm{softmax}(\mathcal{A}^{T})\times\left(\mathcal{L}\mathcal{W}^{v}_{L}\right),$ (3)
where $\mathcal{W}^{k}_{V}$, $\mathcal{W}^{v}_{V}$, $\mathcal{W}^{k}_{L}$, $\mathcal{W}^{v}_{L}$ are learnable weights that map visual and linguistic representations into attention space.
As shown in Fig. 3, after generating the object response in $\mathcal{F}^{V\sim L}$ (textual activation of $H_{1}W_{1}$ pixels) and enhancing the informative tokens in $\mathcal{F}^{L\sim V}$ (visual response for $N$ tokens), we perform high-level cross-modal supplement between them to generate the fine-grained semantic map $\mathcal{S}$:
$\mathcal{S}=\mathcal{F}^{L\sim V}\times(\mathcal{F}^{V\sim L})^{T}.$ (4)
In this manner, the key tokens related to the referent are anticipated to emerge significant responses in the semantic map while suppressing the activation of non-related regions. Considering that spatial relationships between instances and empty-target identification are crucial in GRES, solely relying on the last plain visual feature [20, 11] for mask generation potentially causes the misinterpretation of the referred target. Therefore, we provide an effective way to address the issue, which is the direct utilization of the fine-grained semantic map as input $\mathcal{M}_{1}\in\mathbb{R}^{N\times H_{1}W_{1}}$ to the following Multi-level Mask Aggregation module.
Figure 3: Structure of STM and STQ module. STM takes visual and linguistic features as input and enables bi-directional modality calibration to generate the corresponding fine-grained semantic map at each level. Subsequently, STQ transfers modality complementary contents at the query level with the assistance of the referring semantics and original linguistic information. Further, we visualize the semantic map at each level, showcasing the semantic alignment at different granularities.
Semantic To Query Module.
The textual description in GRES weighs more crucial than the classic RES task in two folds: 1) It additionally encompasses implicit or complex multiple-target information such as counting expressions and relationships among various domains of instances. 2) The empty-target description appears relevant to the image, which will easily result in undesirable object responses. However, previous methods [20, 11] directly utilize the outputs from visual-attended cross attention, i.e., $\mathcal{F}^{L\sim V}$ for the consecutive cross-modal interaction. This approach potentially contributes to the loss of original linguistic information as contents in visual features are intensive but mostly irrelevant.
To address the issues, we propose the Semantic-to-Query (STQ) module, which performs comprehensive multi-modal complementary at the query level. It well incorporates the important referring semantics and preserves valuable language information simultaneously. Formally, we leverage fully connected layers to project the fine-grained semantic map $\mathcal{S}$ into query space:
$\mathcal{Q}^{S}=\mathrm{MLP}(\mathcal{S}).$ (5)
Then, as shown in Fig. 3, the high-level semantics in $\mathcal{Q}^{S}$, instead of the single visual modality, is injected into the language query $\mathcal{Q}^{L}$, i.e., the initial input query $\mathcal{L}^{O}$ or the one from previous STQ module:
$\mathcal{Q}^{L^{\prime}}=\mathrm{MHA}\left(\mathcal{Q}^{L},\mathcal{Q}^{S},\mathcal{Q}^{S}\right),$ (6)
where $\mathrm{MHA}$ is the vanilla multi-head attention mechanism and $\mathcal{Q}^{L^{\prime}}\in\mathbb{R}^{N\times D}$.
As illustrated in Fig. 2, $\mathcal{Q}^{L^{\prime}}$ is further processed in the next-level STM module. It is anticipated that details within visual features become more adequate as the scale increases, which necessitates the more fine-grained modality interaction. Moreover, to prevent the shifted emphasis on original textual information and facilitate a deeper understanding of that with incorporated holistic semantics, we perform language reactivation by reintroducing the original linguistic feature $\mathcal{L}$. Thus, $\mathcal{Q}^{L^{\prime}}$ gets greatly enriched by the progressive modality complementary.
Multi-level Mask Aggregation.
Multiple-target referents typically vary in shape or scale within an image. The existing GRES mask decoding methods which obtain segmentation directly by leveraging the information-redundant joint feature, may lead to undesirable object response in multiple/non-target scenarios. Therefore, we propose Multi-level Mask Aggregation (MMA) to unify the semantic clue progressively for final mask predictions. In MMA, the corresponding object responses of semantic-enriched masks from different levels are hierarchically aggregated, which enhances the activation of proper regions while suppressing the unrelated ones.
Specifically, as shown in Fig. 2, the outputs $\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}\}$ from cascaded STMs are gathered into MMA for final segmentation. Firstly, we convert each level of semantic-enriched segmentation into a two-class mask by multiplying dynamic kernel $B$:
$\mathcal{M}_{i}^{\prime}=\mathcal{M}_{i}\times B_{i},$ (7)
where $B_{i}$ is conditional on $\mathcal{Q}^{L^{\prime}}_{i}$ from each STQ module. Subsequently, they are weighted aggregated through sequential upsampling in a multi-level paradigm, which can be expressed as:
[…]
where $\mathrm{Up}(\cdot)$ indicates bilinear upsampling operation and $\alpha$ controls the fusion ratio between levels.
3.3 Multi-Adaptive Object Counting
Existing GRES methods tend to directly adopt a binary classification branch to judge the existence of objects. However, we observe that such simple design is not completely applicable in general referring segmentation settings. Particularly, it broadly categorizes multiple-target and single-target scenarios into one objective, ignoring their inherent specificity, which can easily incur ambiguity.
Meanwhile, several studies [21, 11] use the language reconstruction technique. Yet, it is not suitable in GRES since reconstructing unrelated content increases the burden of precise object awareness. Therefore, we propose Multi-Adaptive Object Counting (MAC) to endow our model with category-specific counting ability in multiple/single/non-target situations, which facilitates the understanding of both complicated sentences, e.g., counting, compound structure expression, and visual scenes.
Formally, the multi-modal semantic queries $\{\mathcal{Q}^{L^{\prime}}_{1},\mathcal{Q}^{L^{\prime}}_{2},\mathcal{Q}^{L^{\prime}}_{3}\}$ corresponding to each grain are transformed into the counting vector of $C$ categories, respectively:
$\mathcal{C}_{i}=\mathrm{MLP}(\mathcal{Q}^{L^{\prime}}_{i}),$ (9)
where $\mathcal{C}_{i}\in\mathbb{R}^{N\times C}$. Then $\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ are incorporated for the counting prediction $\mathcal{C}_{all}\in\mathbb{R}^{N\times C}$ by average pooling. Finally, we provide the explicit supervision on $\mathcal{C}_{all}$ by computing the $\mathit{L_{1}}$ distance between the multi-level accumulated results $\mathcal{C}^{pred}$ and the ground truth $\mathcal{C}^{gt}$, where $\mathcal{C}^{gt}$ (see in Fig. 2) is annotated as the exact counts of each category. The whole process can be formulated as:
[…]
$\mathit{SmoothL1}(x)$ =$=\begin{cases}0.5\cdot x^{2}&\text{if }|x|<1.\\ |x|-0.5&\text{otherwise}.\end{cases}$
On one hand, the MAC module is designed to seamlessly adapt to the specificity in generalized scenes, i.e., multiple/single/non-target tasks, which benefits the general object matching. On the other hand, guided by the the precise categories and counts of referent targets, our method facilitates comprehensive object perception in both image and language wise.
3.4 Output and Loss
[…]
Therefore, total loss with balance factor $\lambda$ can be formulated as:
[…]
4 Experiment
4.1 Experiment Setups
We evaluate our model on three GRES benchmarks using official evaluation metrics: gRefCOCO [20], R-RefCOCO/+/g [30] and Ref-ZOM [11]. Following previous GRES methods, we also conduct experiments on the RES dataset: RefCOCO/+/g [15, 27]. Elaborate descriptions of each benchmark and metrics are in Sec. A.1, Sec. A.2. We adopt Swin Transformer [25] and BERT [5] as the backbone. Due to space limitations, complete implementation details are in Sec. A.3.
Generalized Referring Expression Segmentation on gRefCOCO
Method Backbone Validation Set Test Set A Test Set B
gIoU cIoU N-acc. gIoU cIoU N-acc. gIoU cIoU N-acc.
MLLM Methods
LISA-V-7B [17] SAM-ViT-H 32.21 38.72 2.71 48.54 52.55 6.37 39.65 44.79 5.00
GSVA-V-7B [32] SAM-ViT-H 63.32 61.70 56.45 70.11 69.23 63.50 61.34 60.26 58.42
LISA-V-7B [17] (ft) SAM-ViT-H 61.63 61.76 54.67 66.27 68.50 50.01 58.84 60.63 51.91
GSVA-V-7B [32] (ft) SAM-ViT-H 66.47 63.29 62.43 71.08 69.93 65.31 62.23 60.47 60.56
Specialist Methods
MattNet [34] ResNet-101 48.24 47.51 41.15 59.30 58.66 44.04 46.14 45.33 41.32
LTS [14] DarkNet-53 52.70 52.30 - 62.64 61.87 - 50.42 49.96 -
VLT [6] DarkNet-53 52.00 52.51 47.17 63.20 62.19 48.74 50.88 50.52 47.82
CRIS [24] CLIP-R101 56.27 55.34 - 63.42 63.82 - 51.79 51.04 -
ReLA [20] Swin-T 56.87 57.73 44.07 65.29 66.02 48.68 57.37 56.86 50.31
HDC (Ours) Swin-T 65.19 62.59 60.32 69.50 68.84 60.44 61.15 59.82 60.07
LAVT [33] Swin-B 58.40 57.64 49.32 65.90 65.32 49.25 55.83 55.04 48.46
ReLA [20] Swin-B 63.60 62.42 56.37 70.03 69.26 59.02 61.02 59.88 58.40
HDC (Ours) Swin-B 68.28 65.42 63.38 72.52 71.60 65.29 63.85 62.79 60.68
Table 1: Comparison with the state-of-the-art methods on gRefCOCO. -V-7B means Vicuna-7B. (ft) denotes that the model is finetuned on the training set of gRefCOCO. denotes that the results are from the official repository.
Method R-RefCOCO R-RefCOCO+ R-RefCOCOg
mIoU mRR rIoU mIoU mRR rIoU mIoU mRR rIoU
CRIS [24] 43.58 76.62 29.01 32.13 72.67 21.42 27.82 74.47 14.60
EFN [7] 58.33 64.64 32.53 37.74 77.12 24.24 32.53 75.33 19.44
VLT [6] 61.66 63.36 34.05 50.15 75.37 34.19 49.67 67.31 31.64
LAVT [33] 69.59 58.25 36.20 56.99 73.45 36.98 59.52 61.60 34.91
LAVT+ [33] 54.70 82.39 40.11 45.99 86.35 39.71 47.22 81.45 35.46
RefSegformer [30] 68.78 73.73 46.08 55.82 81.23 42.14 54.99 71.31 37.65
HDC (Ours) 74.35 83.69 52.81 64.85 87.51 49.09 65.11 84.19 43.85
Table 2: Comparison with state-of-the-art methods on the R-RefCOCO/+/g dataset.
4.2 Main Results
Results on GRES tasks.
To verify our effectiveness in the generalized setting, we first compare our method with previous specialist methods in Table 1 on gRefCOCO [20]. It can be seen that our method achieves new state-of-art performance on all metrics of three evaluation sets in the large-scale GRES benchmark. Under fair comparison with the previous SOTA method ReLA [20], HDC achieves superior performance with a significant margin of $+4.7\%$, $+2.5\%$, $+2.8\%$ in gIoU of val, testA, testB set. Then, we report our results on Ref-ZOM benchmark [11](see in Table 3). As illustrated, our method outperforms all methods under a fair setting, e.g., $+6.3\%$ in Acc., $+1.1\%$ in oIoU. It is worth noting that our method is even better than GSVA [32], which utilizes Multi-Modal Large Language Model (MLLM) [23]. In addition, we further make a comparison on R-RefCOCO/+/g [30]. As depicted in Table 2, our method yields a significant improvement of $+6.7\%$, $+7.0\%$, $+6.2\%$ rIoU in R-RefCOCO/+/g compared with RefSegformer [30].
Results on RES tasks.
Following existing GRES methods [20, 11, 30], we train on each RES dataset without additional data [33]. Due to the main focus on generalized formulation and space limitations, please see the Sec. B.1. HDC also outperforms all GRES methods in RES tasks.
      Method       Backbone       Ref-ZOM Test Set
      oIoU       mIoU       Acc.
      MLLM Methods
      LISA-V-7B [17]       SAM-ViT-H       61.46       60.14       72.58
      GSVA-V-7B [32]       SAM-ViT-H       67.98       67.12       82.66
      LISA-V-7B [17] (ft)       SAM-ViT-H       65.39       66.41       93.39
      GSVA-V-7B [32] (ft)       SAM-ViT-H       68.13       68.29       94.59
      Specialist Methods
      MCN [26]       DarkNet-53       54.70       55.03       75.81
      CMPC [12]       ResNet-101       55.72       56.19       77.01
      VLT [6]       DarkNet-53       60.43       60.21       79.26
      LAVT [33]       Swin-B       64.78       64.45       83.11
      DMMI [11]       Swin-B       68.21       68.77       87.02
      HDC (Ours)       Swin-B       69.31       68.81       93.34
Table 3: Comparison with state-of-the-art methods on the Ref-ZOM dataset.
HSD MAC gIoU cIoU N-acc.
$53.52$ $56.55$ $41.11$
$57.63$ $58.56$ $49.99$
$63.15$ $60.67$ $55.84$
$\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$
Table 4: Effectiveness of the modules.
STM STQ MMA gIoU cIoU N-acc.
$57.63$ $58.56$ $49.99$
$59.70$ $59.95$ $47.59$
$62.06$ $62.36$ $52.30$
$\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$
Table 5: Effects of components of HSD module.
Mask Decode gIoU cIoU
Mask Attn. 63.36 61.97
Deep Sup. 64.11 61.98
MMA $\mathbf{65.19}$ $\mathbf{62.59}$
Both 63.35 62.34
Table 6: Comparison of decoding methods.
MAC Object Indicator gIoU cIoU N-acc. C-acc.
binary head $63.15$ $60.67$ $55.84$
binary head $63.72$ $61.59$ $56.09$ $72.73$
count head $\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$ $\mathbf{74.58}$
Table 7: Impact of different strategies of object-existence identification.
Semantics Lang.Reactive gIoU cIoU N-acc.
60.55 61.19 51.15
63.31 61.56 56.10
65.19 62.59 60.32
Table 8: Impact of different structure of STQ.
4.3 Ablation Study
In this section, we conduct analysis experiments on the gRefCOCO [20] val set using Swin-Tiny as the visual backbone. Note that, we refer to additional ablation studies in Sec. B.2.
Module Effectiveness.
We experiment with the effectiveness of the proposed Hierarchical Semantic Decoding (HSD) and Multi-Adaptive Object Counting (MAC) module. As shown in Table 8, With HSD hierarchically performing modality complementary and multi-level semantic aggregation, the performance is boosted about $4.1\%$ cIoU. With the explicit counting ability, there is an improvement of $2\%$ cIoU. The combination contributes to superior performance on all benchmarks.
Analysis of HSD.
To investigate the validity of three key modules involved in HSD, we provide an analysis of the Semantic-to-Mask (STM), Semantic-to-Query (STQ), and Multi-level Mask Aggregation (MMA) components. As Table 8 shows, the mutual modality complementary for fine-grained semantic map generation at each level helps gain $2.1\%$ gIoU. Then, STQ formulates a higher level understanding of the linguistic information via the complementary of the referring holistic semantics, which brings a clear improvement of $2.3\%$ gIoU. since MMA aggregates across multi-level visual-linguistic semantics, the ablated of it leads to the degradation of $3.1\%$ gIoU.
Variants of Object-Existence Identification.
We investigate the impact of different design variants on the object-existence justification that can be directly reflected by N-acc. As Table 8 shows, if maintaining the binary classification across all situations [20, 30], i.e., MAC is degraded to an auxiliary task, the counting ability still brings benefit. Furthermore, to explicitly quantify the counting ability, we calculate the accuracy between the prediction object count of each category and the true one, termed as C-acc. It can be seen that the simple binary head hampers the object perception, which results in $1.9\%$ C-acc drop. With the subsequent design of counting head upon MAC which facilitates the comprehensive object perception under each situation, HDC yields the most robust performance.
Analysis of STQ Structure.
Results on Table 8 demonstrate that the incorporated referring semantics at the query level is crucial, which contributes to $+2.8\%$ gIoU. Moreover, since the subsequent language reactivation prevents the shifted emphasis of key linguistic contents and formulates a deeper understanding of them via the semantics supplement, the absence of it causes $1.9\%$ gIoU degradation.
Mask Decoding Method.
To prove the effectiveness of MMA, we compare different prevalent mask decoding methods. As shown in Table 8, replacing MMA with mask attention [2] mechanism at each level to generate the final mask prediction is inferior. This can be attributed to the restriction on specific regions at the preliminary level is susceptible to the following modality interaction. Moreover, we observe that the utilization of deep supervision at each level with or without aggregation falls below anticipated outcomes. We speculate that it deteriorates the nature of the hierarchical structure that target responses from levels are different but adaptively reciprocal.
Comprehensive Evaluation in Generalized Settings.
Figure 4: Comprehension evaluation from three perspectives.
To quantify our overall superiority in addressing the challenges introduced in GRES, we synthetically compare HDC with the previous SOTA method, ReLA [20] from three perspectives. The horizontal axis of Fig. 4 denotes the gIoU obtained from only Multiple-target scenarios, the vertical axis indicates the N-acc. calculated in Zero-target cases individually, while the number inside the circle represents the total Pr@0.7 of the One-target situation. As illustrated, HDC shows absolute dominance across dimensions compared to ReLA.
Figure 5: Segmentation Results. (a) and (b) are input images and segmentation results of our HDC with different referring expressions. The term Count specifies the output of MAC module.
Figure 6: Visualization comparison in generalized setting. (a) and (b) are the segmentation results of ReLA and our HDC, respectively. The term Count specifies the output of MAC module.
4.4 Qualitative Results
Benefiting from the hierarchical semantic aggregation with well-aligned granularities and the counting ability, Fig. 5 and Fig. 6 demonstrate our superiority in processing the complex spatial relationship between instances and intricate referring text in multiple/single/non-target scenarios. In contrast, ReLA [20] fails to adequately address these challenges in GRES due to its incomplete modeling of visual-linguistic correlation. More visualization results are in Appendix C.
5 Conclusion
In this paper, we propose a framework called HDC for GRES to boost the multi-grained comprehension of complex referring semantic associations at different levels. With progressively transferring modality complementary information across granularities, HDC aggregates each visual-linguistic correspondence for multi-level semantic decoding. Moreover, with well-aligned semantic contexts, we empower HDC with category-specific counting ability to promote comprehensive object perception in multiple/single/non-target scenarios. Experimental results show the dominance of HDC in all benchmarks and the superiority in addressing the new challenges of GRES.
Appendix
Appendix A Additional Details on Experiment Setup
A.1 Datasets
gRefCOCO.
It contains 278,232 expressions, which includes 80,022 multiple-target referents and 32,202 empty-target ones. There are 60,287 distinct instances being referred in 19,994 images. The images are split into four subsets: training, validation, test-A, and test-B following the same UNC partition of RefCOCO [15].
Ref-ZOM.
Ref-ZOM are selected from COCO dataset [19], which consists of 55,078 images and 74,942 annotated objects. 43,749 images and 58,356 objects are utilized in training, and 11,329 images and 16,586 objects are employed in testing. It is annotated with three different settings, i.e., one-to-zero, one-to-one, one-to-many, each of which corresponds to the empty-target, single-target, and multiple-target in GRES respectively.
R-RefCOCO.
There are three different sets in the dataset, R-RefCOCO, R-RefCOCO+, R-RefCOCOg, and each of which is based on the classic RES benchmark, RefCOCO+/g [15]. Only the validation set follows the UNC partition principle and it is officially stated for evaluation. The formulation rule of the dataset is adding negative sentences into the training set at a 1:1 ratio relative to the positive sentences.
RES.
RefCOCO [15], RefCOCO+ [15], and RefCOCOg [27] are three standard RES benchmarks, each of which contains 19,994, 19,992, and 26,711 images, with 50,000, 49,856, and 54,822 annotated objects and 142,209, 141,564, and 104,560 annotated expressions, respectively.
A.2 Metrics
For GRES, following [20], we measure the effectiveness of our model by Pr@0.7, gIoU, cIoU and N-acc for gRefCOCO. Meantime, oIoU, mIoU, Acc. used in [11] are adopted for Ref-ZOM. In addition, the standard metrics mIoU, mRR, rIoU are for R-RefCOCO [30]. It is important to note that these metrics are officially specified in each respective benchmark.
Similar to the mean IoU, the Generalized IoU (gIoU) computes the average IoU value for each image across all instances. For the empty-target cases, the IoU values for true positive empty-target instances are considered to be $1$ whereas the IoU values for false negative instances are deemed to be 00. cIoU calculates the total intersection pixels over the total union pixels.
mIoU, oIoU is adopted in Ref-ZOM [11], where mIoU is the average IoU value for each image across all cases containing referred objects. oIoU is the same as cIoU. As for R-RefCOCO [30], we use the metric rIoU for quantifying the quality of robust segmentation, which takes the negative sentences into consideration and explicitly assigns the equal weight of the positive one in mIoU calculation. Note that, N-acc. in gRefCOCO, Acc. in Ref-ZOM are in the same formulation where they denote the ratio of the correctly classified empty-target expressions over all the empty-target expressions in the dataset. Similarly, mRR in R-RefCOCO calculates the empty-target expression recognition rate of each image and averages these across the entire dataset.
A.3 Implementation Details
[…]
The model is trained with AdamW optimizer with a weight decay of 0.05. The batch size is set to 48. The learning rate is initialed as 2e-5 and scheduled by cosine learning rate decay by default. Following [33, 20], the input images are resized to 480$\times$480 and the maximum length of referring expressions is set as $20$ for all datasets. Other hyperparameters of the encoding process are the same as ReLA [20]. All experiments are conducted with 8×8$\times$8 × A10 and each takes up about one day with 13GB $\sim$ 18GB memory occupied i.e., it depends on the backbone.
Appendix B Additional Experiments
B.1 Performance on Classic RES Task
As shown in Table 9, our performance is higher than the previous state-of-art GRES methods with a consistent improvement of $1\sim 3\%$ on three datasets, showing that the multi-level semantic aggregation with modality alignment at each granularity and the category-specific object counting capability remains effective in the simpler setting.
Methods Backbone RefCOCO RefCOCO+ RefCOCOg
val test A test B val test A test B val${}_{\text{(U)}}$ test${}_{\text{(U)}}$
MCN [26] Darknet53 62.44 64.20 59.71 50.62 54.99 44.69 49.22 49.40
VLT [6] Darknet53 67.52 70.47 65.24 56.30 60.98 50.08 54.96 57.73
ReSTR [16] ViT-B 67.22 69.30 64.45 55.78 60.44 48.27 - -
CRIS [24] CLIP-R101 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
LAVT [33] Swin-B 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
RefSegformer [30] Swin-B 73.73 75.64 70.09 63.50 68.69 55.44 62.56 63.07
ReLA [20] Swin-B 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
DMMI [11] Swin-B 74.13 77.13 70.16 63.98 69.73 57.03 63.46 64.19
Ours Swin-B 74.81 78.32 70.95 66.80 71.88 58.90 65.59 66.40
Table 9: Comparison with previous state-of-the-art GRES methods on classic RES benchmarks.
$\lambda_{mask}$ $\lambda_{count}$ gIoU cIoU N-acc. C-acc.
$2$ $0.5$ $63.23$ $62.37$ $54.88$ $73.28$
$1$ $0.5$ $62.39$ $61.53$ $54.81$ $73.20$
$1$ $1$ $63.06$ $61.80$ $55.34$ $71.88$
$5$ $0.1$ $64.15$ $62.54$ $57.24$ $71.46$
$2$ $0.1$ $\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$ $\mathbf{74.58}$
Table 10: Results of different ratios of loss.
$\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ gIoU cIoU N-acc. C-acc.
$0.1$ $1.0$ $1.0$ $64.34$ $62.27$ $58.57$ $72.98$
$1.0$ $0.2$ $1.0$ $63.62$ $61.96$ $55.99$ $73.50$
$1.0$ $1.0$ $0.2$ $63.11$ $62.25$ $54.32$ $72.95$
$0.2$ $0.2$ $1.0$ $62.79$ $62.28$ $54.04$ $70.93$
$1.0$ $1.0$ $1.0$ $\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$ $\mathbf{74.58}$
Table 11: Results of different aggregation degrees among levels in MMA.
B.2 Additional Ablation Studies
Loss Ratio.
[…]
Different Aggregation Degrees.
The $\alpha_{i}$ controls the fusion ratio between levels in Multi-level Mask Aggregation module, which represents the importance of each level in the final mask generation. Therefore, We investigate the effect of it in Table 11. As illustrated, ignorance of any level will lead to inferior performance. We speculate that the complex spatial relationship between instances with the intricate referring expression highlights the benefits of hierarchical structures where information in each level is adaptively reciprocal.
Appendix C Visualizations.
Figure 7: Visualization in Multi-Level Mask Aggregation. (a) and (c) indicate the input image and corresponding segmentation result of our HDC, respectively. (b) illustrates the aggregated activation map at each level.
C.1 Multi-level Mask Aggregation Visualization.
As illustrated in the main paper, we hierarchically aggregate each visual-linguistic correspondence in different granularity for multi-level semantic decoding. To further prove the rationality of our design, we visualize the aggregated activation map (originally from the semantic map) at each level in Fig. 7. As observed, with progressive mutual modality complementary across granularities, the desired regions of the referent target are fully activated with the increase of the levels.
C.2 Additional Segmentation Comparisons.
As illustrated in the main paper and Fig. 6, HDC can better handle the extreme challenges of GRES under multiple/single/empty target scenarios compared with the previous SOTA method ReLA [20]. Here, we incorporate more cases in Fig. 9 to demonstrate our superiority in modeling the intricate referring association between visual and linguistic.
C.3 Segmentation Results of HDC.
Fig. 8 demonstrates the effectiveness of HDC in complex generalized settings, e.g., complex geometrical relationships between instances, deceptive objects, compound sentence structures, and intricate associations between referring expressions and images.
Appendix D Limitations.
Benefiting from the hierarchical semantic decoding and the explicit counting ability, our method HDC sufficiently addresses the limitations of existing GRES and achieves superiority in meeting the challenges of GRES. However, there are some potential limitations. Since the referring sentences in GRES contain multiple-target expressions, the truncation of the input text into $20$ may lose detailed descriptions of some of the targets. Although we compensate it with sentence-level textual features, it still remains unsatisfactory in terms of the incompleteness of fine-grained target description. We think that how to fully utilize the textual expression in GRES is an interesting future direction.
Appendix E Error Bar.
We have retrained our model several times on gRefCOCO [20]. The results demonstrate that the randomness of the model has little effect on the performance, i.e., the standard variation is about $0.36$.
Appendix F Broader Impacts.
Malicious use of the GRES model may lead to potential negative societal impacts, including but not limited to unauthorized surveillance or privacy-infringing tracking. However, we firmly believe that the task itself is neutral with positive implications, such as image editing and human-robot interaction.
Figure 8: Segmentation results of our method in generalized settings. (a) denotes the input image, and (b) showcases the segmentation results of HDC under multiple/single/non-target situations with different referring expressions.
Figure 9: Segmentation results comparison in the generalized setting. (a) denotes the input image, (b) and (c) are segmentation results of ReLA and our HDC, respectively. Empty, Single, and Multiple target referent situations with different referring expressions are in top-to-bottom accordance.

### v2

@@
-HDC: Hierarchical Semantic Decoding with Counting Assistance for Generalized Referring Expression Segmentation
+CoHD: A Counting-Aware Hierarchical Decoding Framework for Generalized Referring Expression Segmentation
 Abstract
-The newly proposed Generalized Referring Expression Segmentation (GRES) amplifies the formulation of classic RES by involving multiple/non-target scenarios.
-Recent approaches focus on optimizing the last modality-fused feature which is directly utilized for segmentation and object-existence identification.
-However, the attempt to integrate all-grained information into a single joint representation is impractical in GRES due to the increased complexity of the spatial relationships among instances and deceptive text descriptions.
-Furthermore, the subsequent binary target justification across all referent scenarios fails to specify their inherent differences, leading to ambiguity in object understanding.
-To address the weakness, we propose a Hierarchical Semantic Decoding with Counting Assistance framework (HDC).
-It hierarchically transfers complementary modality information across granularities, and then aggregates each well-aligned semantic correspondence for multi-level decoding.
-Moreover, with complete semantic context modeling, we endow HDC with explicit counting capability to facilitate comprehensive object perception in multiple/single/non-target settings.
-Experimental results on gRefCOCO, Ref-ZOM, R-RefCOCO, and RefCOCO benchmarks demonstrate the effectiveness and rationality of HDC which outperforms the state-of-the-art GRES methods by a remarkable margin.
-Code will be available here.
+The newly proposed Generalized Referring Expression Segmentation (GRES) amplifies the formulation of classic RES by involving complex multiple/non-target scenarios.
+Recent approaches address GRES by directly extending the well-adopted RES frameworks with object-existence identification.
+However, these approaches tend to encode multi-granularity object information into a single representation, which makes it difficult to precisely represent comprehensive objects of different granularity.
+Moreover, the simple binary object-existence identification across all referent scenarios fails to specify their inherent differences, incurring ambiguity in object understanding.
+To tackle the above issues, we propose a Counting-Aware Hierarchical Decoding framework (CoHD) for GRES.
+By decoupling the intricate referring semantics into different granularity with a visual-linguistic hierarchy, and dynamic aggregating it with intra- and inter-selection, CoHD boosts multi-granularity comprehension with the reciprocal benefit of the hierarchical nature.
+Furthermore, we incorporate the counting ability by embodying multiple/single/non-target scenarios into count- and category-level supervision, facilitating comprehensive object perception.
+Experimental results on gRefCOCO, Ref-ZOM, R-RefCOCO, and RefCOCO benchmarks demonstrate the effectiveness and rationality of CoHD which outperforms state-of-the-art GRES methods by a remarkable margin.
+Code is available at here.
+00footnotetext: 🖂 Corresponding author.
 1 Introduction
+Figure 1: Comparison of decoding paradigms.
+Previous GRES methods attempt to integrate multi-granularity context into one joint feature for mask and object-existence prediction (a), which exhibits vulnerability to the complexity of multiple/non-target scenarios (c).
+Instead of relying on the single representation, we decouple the referring correspondence into the semantic hierarchy of different granularity and perform dynamic selective aggregation for robust segmentation.
+Further, in contrast to simple binary classification, we empower CoHD with counting ability to promote object perception (b).
 Referring Expression Segmentation (RES) is one of the fundamental tasks in visual-linguistic understanding, which aims at localizing the referred object in an image with a natural language description.
@@
 However, it poses new challenges with the additional generalized settings.
-The multiple-target referent necessitates more precise semantic clues to handle complicated spatial-text relationships between instances for arbitrary location segmentation.
+The multiple-target referent necessitates more comprehensive semantic clues to handle complicated spatial-text relationships between instances for arbitrary location segmentation.
 On the other hand, the empty-target one highlights the significance of contextual information in mitigating undesirable responses from similar objects.
-Therefore, an intuitive approach is to endow prevalent RES paradigms [33, 6, 7] with new capability.
+An intuitive approach is to endow prevalent RES methods [36, 5, 6] with multi/non-target referring capability.
 For instance, ReLA [22] introduces a large-scale generalized dataset gRefCOCO and utilizes spatial-aware queries to transform GRES into a related region retrieval task.
 Meanwhile, DMMI [10] and RefSegformer [33] enhance modality interaction in the early-fusion process and incorporate new benchmarks Ref-ZOM and R-RefCOCO respectively, to realize robust segmentation.
-Figure 1: Illustration of different decoding processes.
-Previous GRES methods attempt to integrate all-grained context into one joint feature which is directly for mask and object-existence prediction (a), exhibiting vulnerability to challenges of multiple/non-target scenarios (c).
-Instead of relying on the single representation, we perform hierarchical semantic decoding with multi-level visual-linguistic correspondence aggregation for robust segmentation.
-Moreover, in contrast to simple binary classification, we empower HDC with counting ability to promote overall object awareness (b).
-Although these methods have made advancements in GRES, they still have intrinsic weaknesses in two aspects: 1) They are limited by the bottleneck of capturing complete associations between visual and linguistic contents.
-As Fig.
-1 shows, they strive to optimize the last modality-fused feature which is directly for final mask prediction.
-However, the extra complexity of text descriptions and geometric relationships among instances in GRES impedes the integration of all-grained contexts into one joint representation.
-Without complete modality alignment at each granularity, it potentially leads to incorrect or incomplete object response.
-2) A simple binary classification head for object-existence identification not only occurs vulnerability to deceptive instances with accumulated bias in single representation but also incurs ambiguity for ignoring inherent differences of one-one/many referents.
-Therefore, the significant problem lies in how to excavate the potential of multi-grained semantic correspondence between visual and linguistic when decoding, as well as fully facilitate the comprehensive object perception by considering the specificity of each scenario.
-Motivated by that, we present a Hierarchical Semantic Decoding with Counting Assistance framework (HDC).
-In HDC, a Hierarchical Semantic Decoding module (HSD) progressively transfers complementary information between modalities across different granularities to formulate intricate instance-linguist correlation hierarchies.
-Instead of leveraging single joint features, HSD utilizes multi-level semantic aggregation for robust segmentation.
-With hierarchical semantic-context alignment, we further introduce Multi-Adaptive Object Counting module (MAC) on HSD to empower the model with explicit counting ability, which improves overall object awareness under each referring setting.
-We conduct extensive experiments on three popular GRES benchmarks, i.e., gRefCOCO [20], Ref-ZOM [11] and R-RefCOCO [30] and RES datasets RefCOCO/+/g [15, 27] to validate the effectiveness of our methods.
-Results show that HDC sufficiently meets the challenges of GRES and addresses the limitations of existing methods, which notably outperforms them for all benchmarks.
+Despite advancements in GRES, the direct extension from RES to GRES exhibits less compatibility in two aspects: 1) Incomplete Semantic Context: as shown in Fig.
+1 (a) and (c), the previous decoding paradigm is solely dependent on a single representation for multi-granularity modeling, which leads to incomplete or incorrect object responses.
+We speculate that primarily stems from the increased complexity of referring semantics in GRES, e.g., intricate spatial relationships among instances, and deceptive expressions.
+It easily incurs the shifted emphasis on both visual and language-wise, thereby impeding the integration of all-grained contexts into one joint representation.
+2) Imprecise Object Perception: previous GRES methods adopt simple binary classification for object-existence identification, which ignores the inherent distinction between multiple and single referent scenarios, potentially incurring ambiguity in object understanding.
+To well address the challenges of GRES and break through the limitations of previous methods, we delve into a further investigation.
+Firstly, the extra referring scenarios, i.e., multiple/empty-target referents, bring more difficulty in accurately activating complete regions and discerning the non-related ones.
+For instance, as shown in Fig.
+1 (b), different referring object activations vary in shapes, scales, and locations, and misinterpretation potentially occurs at the latter level.
+It indicates that simply relying on a single joint feature for mask decoding is insufficient.
+Instead, the complex referring semantics should be decoupled by different granularity and aggregated hierarchically for robust segmentation.
+Secondly, the inherent distinction of multiple/single referent cases should not be ignored since it makes a difference to the precision object perception when encountering implicit expressions such as “A line of cars”.
+Inspired by that, we present a Counting-aware Hierarchical Decoding framework (CoHD).
+Specifically, a multi-grained visual-linguistic hierarchy is formulated by the Hierarchical Semantic Decoder (HSD).
+Instead of simply leveraging a single joint feature, HSD performs dynamic hierarchical aggregation with intra- and inter-selection.
+Thus, the intricate referring semantic clues can be completely leveraged.
+Furthermore, we introduce the Adaptive Object Counting module (AOC) to specify the inherent distinction among multiple/single/non-target scenarios by embodying each into count- and category-level supervision, which tackles ambiguity and facilitates precise object perception.
+We conduct extensive experiments on three popular GRES benchmarks, i.e., gRefCOCO [22], Ref-ZOM [10] and R-RefCOCO [33] along with RES datasets RefCOCO/+/g [15, 29].
+Experimental results demonstrate that CoHD remarkably outperforms previous works on all benchmarks, showcasing the effectiveness of the proposed method for complicated scenarios in GRES.
 Overall, our contributions are summarized as follows:
-  • We present a Hierarchical Semantic Decoding with Counting Assistance framework (HDC) for GRES to boost the multi-grained comprehension of complex referring semantic associations at different levels.
-In HDC, a Hierarchical Semantic Decoding module (HSD) is designed to progressively transfer mutual modality complementary information across granularities, and utilize each well-aligned semantic correspondence for multi-level decoding.
-  • Built on HSD, we further introduce Multi-Adaptive Object Counting module (MAC) to facilitate comprehensive object perception by assigning explicit counting capability, which seamlessly adapts to the specificity of multiple/single/non-target scenarios.
-To our best knowledge, we are the first to bring object counting into GRES.
-  • Without bells and whistles, our approach outperforms all existing state-of-the-art methods by a remarkable margin under fair comparison, e.g., $+4.7\%$ gIoU in gRefCOCO, $+6.7\%$ rIoU in R-RefCOCO and $+5.3\%$ Acc.
+  • We present the Counting-Aware Hierarchical Decoding framework for GRES, aiming to decouple the complex referring semantics into different levels for boosting multi-granularity comprehension.
+In CoHD, the Hierarchical Semantic Decoder (HSD) formulates a visual-linguistic hierarchy of different granularity and performs mask decoding by dynamic hierarchical aggregation.
+  • We introduce the Adaptive Object Counting module (AOC) to facilitate precise object perception by assigning explicit counting capability, which seamlessly adapts to the specificity of multiple/single/non-target scenarios.
+  • Our approach outperforms all existing state-of-the-art methods by a remarkable margin without bells and whistles, e.g., $+4.8\%$ gIoU in gRefCOCO, $+7.5\%$ rIoU in R-RefCOCO and $+6.3\%$ Acc.
 in Ref-ZOM.
 2 Related Work
+Figure 2: Overview of CoHD.
+The model takes an image with the referring expression as input.
+After the encoding process, the Hierarchical Semantic Decoder (HSD) decouples the referring semantics into different granularity for establishing visual-linguistic hierarchy and then performs dynamic aggregation with intra- and inter-selection for final segmentation.
+Moreover, with well-aligned semantic contexts, we introduce the Adaptive Object Counting module (AOC) to fully promote comprehensive object perception with the category-specific counting ability under multiple/single/non-target scenarios.
 Referring Expression Segmentation.
@@
 To overcome the weakness of insufficiency in modeling visual-text relation that prior works suffer, recent studies [36, 5, 23, 26, 16] design advanced attention-based mechanism [31] to facilitate multi-modal interaction.
-Since previous works are restricted by pre-defined rules that one sentence must match an object, ReLA [20] proposes the Generalized Referring Expression Segmentation (GRES) task, which additionally involves both empty-target and multiple-target scenarios.
-Moreover, DMMI [11] introduces a new benchmark and baseline to realize beyond one target segmentation.
-Similarly, RefSegformer [30] endows the transformer-based model with empty-target sentence discrimination, achieving robust segmentation.
-However, they seek to integrate all-grained contents into a single joint representation for mask prediction, which is impractical in GRES.
-In contrast, we aggregate visual-linguistic semantic clues in hierarchical structures for more generalized segmentation.
+Since previous works are restricted by pre-defined rules that one sentence must match an object, ReLA [22] proposes the Generalized Referring Expression Segmentation (GRES) task, which additionally involves empty-target and multiple-target scenarios.
+DMMI [10] introduces a new benchmark and baseline to realize beyond one target segmentation.
+Similarly, RefSegformer [33] endows the transformer-based model with empty-target sentence discrimination.
+Further, GSVA [35] reuses the [SEG] token and additionally incorporates the [REJ] one into LISA [18], which equips the MLLM with GRES ability.
+However, these methods seek to encode multi-granularity object information into a single joint representation for mask prediction, which is impractical in GRES.
+In contrast, we aggregate visual-linguistic semantic hierarchy of different granularity in a hierarchical manner for more generalized segmentation.
 Object Counting.
@@
 The detection-based methods [12, 7, 8] obtain the number of instances by proposals localization and ranking while the regression-based ones [2, 19, 30] generates predicted counts via density map integration.
-Different from the classic object counting mechanism, our proposed Multi-Adaptive Object Counting module performs at the query level.
+Different from the classic object counting mechanism, our proposed Adaptive Object Counting module performs at the query level.
 By considering various scenarios, i.e., multiple/single/non-target referents, our model achieves comprehensive object perception under each situation.
@@
 3.1 Overview
-Figure 2: Overview of HDC.
-The model takes an image with the referring expression as input.
-After the encoding process, the Hierarchical Semantic Decoding module (HSD) transfers mutual modality complementary information hierarchically for multi-grained alignment and then aggregates each to perform multi-level decoding for final segmentation.
-Moreover, with well-aligned semantic contexts, we introduce Multi-Adaptive Object Counting module (MAC) to fully promote comprehensive object perception by the category-specific counting ability in multiple/single/non-target scenarios.
+Figure 3: Illustration of Hierarchical Semantic Decoder.
+The Semantic Decoding Module takes visual and linguistic features as input and enables bi-directional modality calibration to generate the corresponding fine-grained semantic map at each level, which formulates a visual-linguist hierarchy.
+Subsequently, we incorporate the referring semantics into the query level with the original language information, which facilitates a more holistic object understanding.
+With the well-aligned semantic hierarchy, we perform dynamic aggregation with the intra- and inter-selection mechanism for mask decoding to excavate the potential of reciprocal benefits brought by the hierarchical nature.
 Fig.
-2 depicts the overall pipeline of our HDC.
-Given an image $\mathcal{I}$, the visual backbone, e.g., Swin-Base [25] is to extract a set of vision features $\mathcal{V}_{i}\in\mathrm{R}^{D^{V}\times H_{i}\times W_{i}}$, $i\in\{1,2,3,4\}$, where $H_{i}$ and $W_{i}$ denote the width and height of each scale of feature maps respectively.
+2 depicts the overall pipeline of CoHD.
+Given an image $\mathcal{I}$, the visual backbone, e.g., Swin-Base [27] is to extract a set of vision features $\mathcal{V}_{i}\in\mathbb{R}^{D^{V}\times H_{i}\times W_{i}}$, $i\in\{1,2,3,4\}$, where $H_{i}$ and $W_{i}$ denote the width and height of each scale of feature maps respectively.
 $D^{V}$ signifies the channel dimension.
-$\mathcal{V}_{i}$ is further encoded as mask features $\mathcal{V}^{m}_{i}$ by deformable attention [35].
+$\mathcal{V}_{i}$ is further encoded as mask features $\mathcal{V}^{m}_{i}$ ($i\in\{1,2,3\}$) by deformable attention [38].
 Simultaneously, the textual expression $\mathcal{E}$ with $T$ words is sent into a transformer-based [4] language encoder which generates the word-level embedding $\mathcal{L}^{w}\in\mathbb{R}^{T\times D^{L}}$ and sentence-level embedding $\mathcal{L}^{s}\in\mathbb{R}^{D^{L}}$.
 Considering that $\mathcal{L}^{w}$ contains fine-grained descriptions while $\mathcal{L}^{s}$ expresses the general information, language query $\mathcal{L}^{O}\in\mathbb{R}^{N\times D^{L}}$ is obtained via the concatenation of them.
-Meantime, $\mathcal{L}^{O}$ and $\mathcal{V}^{m}_{i}$ are sent into the Hierarchical Semantic Decoding (HSD) module, which generates corresponding object segmentation $\mathcal{M}\in\mathbb{R}^{H\times W}$ by the progressively mutual modality complementary across multi-granularity and multi-level semantic aggregation.
-In addition, Multi-Adaptive Object Counting (MAC) is designed to promote the precise object perception of different scenarios by category-specific counting ability.
-We elaborate on each module in the following sections.
-3.2 Hierarchical Semantic Decoding Module
-The extended formulation of multiple-target and empty-target referents in GRES brings significant challenges since the former necessities a more precise cross-modal understanding of the complex association between objects, e.g., different domains, geometrical relationships while the latter highlights long-range contextual information for mitigating interference from similar deceptive objects.
-Previous GRES methods [20, 11, 30] struggle to optimize the last modality-fused features for the subsequent mask prediction, ignoring the significance of explicit modality alignment at each granularity.
-While it is acceptable in classic RES, the extremely intricate referring semantics between instances and text in GRES hinder the integration of all-grained modality correspondence in one representation, thus incurring misinterpretation.
-To address the issues, we propose a Hierarchical Semantic Decoding module to perform mutual modality calibration at each granularity and multi-level semantic integration for robust segmentation.
-Specifically, we utilize cascaded Semantic To Mask/Query modules to progressively transfer complementary information between modalities, which captures the high-level visual-linguistic correlation at each hierarchy.
-Then the subsequent Multi-level Mask Aggregation is to perform semantic aggregation for object-aware segmentation.
-Semantic To Mask Module.
-The core idea of STM is to perform bi-directional modality complementary to generate the fine-grained semantic map at each scale of visual features, which is further utilized in the multi-level decoding process.
-As Fig.
-3 shows, the visual feature $\mathcal{V}^{m}_{i}$ and the language query $\mathcal{L}^{O}$ are concurrently input in STM.
-To be specific, taking $\mathcal{V}^{m}_{1}$ as an example, we firstly leverage fully connected layers to map visual features and language query into a shared joint space $D$, respectively.
+$\mathcal{L}^{O}$ and $\mathcal{V}^{m}_{i}$ are simultaneously sent into the Hierarchical Semantic Decoder (HSD), which generates corresponding object segmentation $\mathcal{M}\in\mathbb{R}^{H\times W}$ by dynamic hierarchical mask decoding on the semantic hierarchy of different granularity.
+Subsequently, Adaptive Object Counting (AOC) is designed to promote the precise object perception of different scenarios by counting ability.
+3.2 Hierarchical Semantic Decoder
+Previous GRES methods [22, 10, 33] adopt the traditional RES decoding paradigm which only incorporates the last modality-fused features.
+It is acceptable in RES since it is to match the single referred object through multi-modal interaction.
+However, it is less compatible with GRES.
+We observe that misinterpretation of region activation potentially occurs at the specific granularity in multiple-target cases, which indicates that simply performing sequential matching within each level as RES will lead to incomplete object segmentation.
+For the non-target referent, the final joint representation of multi-grained information is highly redundant, possibly resulting in incorrect object responses.
+Therefore, the hierarchical aggregation across granularity is significant for GRES, since it benefits from the reciprocal properties of semantic activation at each granularity, which helps to address the previous issues.
+Inspired by that, we propose a Hierarchical Semantic Decoder to formulate a visual-linguistic hierarchy through concatenated semantic decoding modules and perform hierarchical dynamic aggregation for robust segmentation.
+Semantic Decoding Module.
+The core idea of SDM is in two folds: 1) Mask Generation: generate the fine-grained semantic map at each scale of visual features which is further utilized in the dynamic aggregation process.
+2) Query Refinement: the well-supervised semantic map helps guide the formulation of holistic comprehension of the referring semantics at the query level.
+Thus, the intricate semantics are well modeled by the mutual-enhanced relationship within modules.
+Specifically, taking $\mathcal{V}^{m}_{1}$ as an example, we firstly leverage fully connected layers to map visual features and language query into a shared joint space $D$, respectively.
 The mapped representations are further denoted as $\mathcal{V}$ and $\mathcal{L}$ for clarity.
-Then, we derive the coarse semantic map $\mathcal{A}$ by attention mechanism and it is applied in visual-linguistic interaction to align information flow between modalities.
-The preliminary modality complementary highlights the corresponding object activation of language-guided visual features, while simultaneously relieving the potential noise in visual-oriented textual representations.
+Then, we derive the coarse semantic map $\mathcal{A}$ by attention mechanism and it is applied in visual-linguistic interaction to align information flow between modalities for generating the fine-grained semantic map $\mathcal{S}$.
+The modality complementary highlights the corresponding object activation of language-guided visual features, while simultaneously relieving the potential noise in visual-oriented textual representations.
 The above process can be formulated as:
@@
 $\mathcal{F}^{V\sim L}=\mathrm{softmax}(\mathcal{A}^{T})\times\left(\mathcal{L}\mathcal{W}^{v}_{L}\right),$ (3)
+$\mathcal{S}=\mathcal{F}^{L\sim V}\times(\mathcal{F}^{V\sim L})^{T},$ (4)
 where $\mathcal{W}^{k}_{V}$, $\mathcal{W}^{v}_{V}$, $\mathcal{W}^{k}_{L}$, $\mathcal{W}^{v}_{L}$ are learnable weights that map visual and linguistic representations into attention space.
-As shown in Fig.
-3, after generating the object response in $\mathcal{F}^{V\sim L}$ (textual activation of $H_{1}W_{1}$ pixels) and enhancing the informative tokens in $\mathcal{F}^{L\sim V}$ (visual response for $N$ tokens), we perform high-level cross-modal supplement between them to generate the fine-grained semantic map $\mathcal{S}$:
-$\mathcal{S}=\mathcal{F}^{L\sim V}\times(\mathcal{F}^{V\sim L})^{T}.$ (4)
-In this manner, the key tokens related to the referent are anticipated to emerge significant responses in the semantic map while suppressing the activation of non-related regions.
-Considering that spatial relationships between instances and empty-target identification are crucial in GRES, solely relying on the last plain visual feature [20, 11] for mask generation potentially causes the misinterpretation of the referred target.
-Therefore, we provide an effective way to address the issue, which is the direct utilization of the fine-grained semantic map as input $\mathcal{M}_{1}\in\mathbb{R}^{N\times H_{1}W_{1}}$ to the following Multi-level Mask Aggregation module.
-Figure 3: Structure of STM and STQ module.
-STM takes visual and linguistic features as input and enables bi-directional modality calibration to generate the corresponding fine-grained semantic map at each level.
-Subsequently, STQ transfers modality complementary contents at the query level with the assistance of the referring semantics and original linguistic information.
-Further, we visualize the semantic map at each level, showcasing the semantic alignment at different granularities.
-Semantic To Query Module.
-The textual description in GRES weighs more crucial than the classic RES task in two folds: 1) It additionally encompasses implicit or complex multiple-target information such as counting expressions and relationships among various domains of instances.
-2) The empty-target description appears relevant to the image, which will easily result in undesirable object responses.
-However, previous methods [20, 11] directly utilize the outputs from visual-attended cross attention, i.e., $\mathcal{F}^{L\sim V}$ for the consecutive cross-modal interaction.
-This approach potentially contributes to the loss of original linguistic information as contents in visual features are intensive but mostly irrelevant.
-To address the issues, we propose the Semantic-to-Query (STQ) module, which performs comprehensive multi-modal complementary at the query level.
+In this manner, the key tokens related to the referent are anticipated to emerge significant responses in the semantic map $\mathcal{S}$ while suppressing the activation of non-related regions.
+Then, it is further utilized as the input $\mathcal{M}_{1}\in\mathbb{R}^{N\times H_{1}W_{1}}$ to formulate the multi-granularity visual-linguistic correspondence hierarchy for dynamic aggregation.
+Considering that the textual description in GRES weighs more crucial than the classic RES task since it encompasses additionally implicit multiple-target information or deceptive description, we further introduce the well-aligned semantics into the query level for better object perception (see in Fig.
+3).
 It well incorporates the important referring semantics and preserves valuable language information simultaneously.
@@
 $\mathcal{Q}^{S}=\mathrm{MLP}(\mathcal{S}).$ (5)
-Then, as shown in Fig.
-3, the high-level semantics in $\mathcal{Q}^{S}$, instead of the single visual modality, is injected into the language query $\mathcal{Q}^{L}$, i.e., the initial input query $\mathcal{L}^{O}$ or the one from previous STQ module:
+The high-level semantics in $\mathcal{Q}^{S}$, instead of the single visual modality, is injected into the language query $\mathcal{Q}^{L}$, i.e., the initial input query $\mathcal{L}^{O}$ or the one from previous SD module:
 $\mathcal{Q}^{L^{\prime}}=\mathrm{MHA}\left(\mathcal{Q}^{L},\mathcal{Q}^{S},\mathcal{Q}^{S}\right),$ (6)
 where $\mathrm{MHA}$ is the vanilla multi-head attention mechanism and $\mathcal{Q}^{L^{\prime}}\in\mathbb{R}^{N\times D}$.
-As illustrated in Fig.
-2, $\mathcal{Q}^{L^{\prime}}$ is further processed in the next-level STM module.
-It is anticipated that details within visual features become more adequate as the scale increases, which necessitates the more fine-grained modality interaction.
-Moreover, to prevent the shifted emphasis on original textual information and facilitate a deeper understanding of that with incorporated holistic semantics, we perform language reactivation by reintroducing the original linguistic feature $\mathcal{L}$.
-Thus, $\mathcal{Q}^{L^{\prime}}$ gets greatly enriched by the progressive modality complementary.
-Multi-level Mask Aggregation.
-Multiple-target referents typically vary in shape or scale within an image.
-The existing GRES mask decoding methods which obtain segmentation directly by leveraging the information-redundant joint feature, may lead to undesirable object response in multiple/non-target scenarios.
-Therefore, we propose Multi-level Mask Aggregation (MMA) to unify the semantic clue progressively for final mask predictions.
-In MMA, the corresponding object responses of semantic-enriched masks from different levels are hierarchically aggregated, which enhances the activation of proper regions while suppressing the unrelated ones.
-Specifically, as shown in Fig.
-2, the outputs $\{\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}\}$ from cascaded STMs are gathered into MMA for final segmentation.
-Firstly, we convert each level of semantic-enriched segmentation into a two-class mask by multiplying dynamic kernel $B$:
-$\mathcal{M}_{i}^{\prime}=\mathcal{M}_{i}\times B_{i},$ (7)
-where $B_{i}$ is conditional on $\mathcal{Q}^{L^{\prime}}_{i}$ from each STQ module.
-Subsequently, they are weighted aggregated through sequential upsampling in a multi-level paradigm, which can be expressed as:
+With greatly enriched by the referring semantics from $\mathcal{Q}^{S}$, we further perform language reactivation by reintroducing the original linguistic feature $\mathcal{L}$, which not only prevents the shifted emphasis but also facilitates a more holistic understanding of that.
+Dynamic Hierarchical Aggregation.
+We observe that the desired object responses typically vary in granularity when performing multiple object referring.
+Similar to the non-target one, the unrelated region activation will potentially occur due to the deceptive descriptions.
+Therefore, we emphasize the importance of reciprocal properties of multi-granularity information in GRES by proposing Dynamic Hierarchical Aggregation to unify the visual-linguistic hierarchy with intra- and inter-selective mechanisms.
+Shown in Fig.
+3, the output $\{\left(\mathcal{M}_{1},\mathcal{Q}^{L^{\prime}}_{1}\right),\left(\mathcal{M}_{2},\mathcal{Q}_{2}^{L^{\prime}}\right),\left(\mathcal{M}_{3},\mathcal{Q}_{3}^{L^{\prime}}\right)\}$ from cascaded SDMs are gathered into DHA for segmentation.
+It is anticipated that $\mathcal{M}_{1}$ with the largest downsampling rate contains the highest level of object semantics, e.g., location, being referred or not, etc.
+In contrast, $\mathcal{M}_{3}$ mainly perceives the shape, and texture character of the objects.
+However, simply aggregation by summation easily introduces bias accumulation, leading to unsatisfactory outcomes.
+To fully excavate the benefit of hierarchical nature, we perform intra- and inter-selection on $\mathcal{M}$ and $\mathcal{Q}^{L^{\prime}}$ respectively.
+1) Inter-Selection: inspired by the current success of MoE [13], we design a gating network to assign the saliency score for each level, which adaptively highlights the related object responses through granularity:
+$\alpha_{i}=\mathcal{G}(\mathcal{Q}_{i}^{L^{\prime}}),i\in\{1,2,3\}$ (7)
+$\mathcal{G}(x):=\mathrm{Sigmoid}(x\mathcal{W}_{g}).$ (8)
+where $\alpha$in $\mathbb{R}^{1}$ and $\mathcal{W}_{g}$ is the learnable weights.
+2) Intra-Selection: affected by the deceptive or compound structure textual description, some tokens may exhibit sparse and cluttered information pertaining to targets in $\mathcal{M}$.
+We assume that the informative and highly related region activation plays a significant role in hierarchical aggregation.
+With the purpose of dynamically enhancing the referred object responses and simultaneously suppressing the non-related ones, we leverage channel attention module [32] on $\mathcal{M}_{i}$ to evaluate the degree of referring correspondence in token level:
+$\mathcal{M}_{i}^{\prime}=\mathrm{ChannelAttn}(\mathcal{M}_{i})\cdot\mathcal{M}_{i}.$ (9)
+After suppressing the unrelated activation based on the low referring correspondence, $\{\mathcal{M}_{1}^{\prime},\mathcal{M}_{2}^{\prime},\mathcal{M}_{3}^{\prime}\}$ are weighted aggregated with each corresponding saliency score $\alpha_{i}$ through sequential upsampling in a multi-level paradigm.
+The whole process can be expressed as:
 […]
-where $\mathrm{Up}(\cdot)$ indicates bilinear upsampling operation and $\alpha$ controls the fusion ratio between levels.
-3.3 Multi-Adaptive Object Counting
-Existing GRES methods tend to directly adopt a binary classification branch to judge the existence of objects.
-However, we observe that such simple design is not completely applicable in general referring segmentation settings.
-Particularly, it broadly categorizes multiple-target and single-target scenarios into one objective, ignoring their inherent specificity, which can easily incur ambiguity.
-Meanwhile, several studies [21, 11] use the language reconstruction technique.
-Yet, it is not suitable in GRES since reconstructing unrelated content increases the burden of precise object awareness.
-Therefore, we propose Multi-Adaptive Object Counting (MAC) to endow our model with category-specific counting ability in multiple/single/non-target situations, which facilitates the understanding of both complicated sentences, e.g., counting, compound structure expression, and visual scenes.
-Formally, the multi-modal semantic queries $\{\mathcal{Q}^{L^{\prime}}_{1},\mathcal{Q}^{L^{\prime}}_{2},\mathcal{Q}^{L^{\prime}}_{3}\}$ corresponding to each grain are transformed into the counting vector of $C$ categories, respectively:
-$\mathcal{C}_{i}=\mathrm{MLP}(\mathcal{Q}^{L^{\prime}}_{i}),$ (9)
-where $\mathcal{C}_{i}\in\mathbb{R}^{N\times C}$.
-Then $\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ are incorporated for the counting prediction $\mathcal{C}_{all}\in\mathbb{R}^{N\times C}$ by average pooling.
-Finally, we provide the explicit supervision on $\mathcal{C}_{all}$ by computing the $\mathit{L_{1}}$ distance between the multi-level accumulated results $\mathcal{C}^{pred}$ and the ground truth $\mathcal{C}^{gt}$, where $\mathcal{C}^{gt}$ (see in Fig.
-2) is annotated as the exact counts of each category.
-The whole process can be formulated as:
-[…]
-$\mathit{SmoothL1}(x)$ =$=\begin{cases}0.5\cdot x^{2}&\text{if }|x|<1.\\ |x|-0.5&\text{otherwise}.\end{cases}$
-On one hand, the MAC module is designed to seamlessly adapt to the specificity in generalized scenes, i.e., multiple/single/non-target tasks, which benefits the general object matching.
-On the other hand, guided by the the precise categories and counts of referent targets, our method facilitates comprehensive object perception in both image and language wise.
-3.4 Output and Loss
-[…]
-[…]
-Therefore, total loss with balance factor $\lambda$ can be formulated as:
-[…]
-4 Experiment
-4.1 Experiment Setups
-We evaluate our model on three GRES benchmarks using official evaluation metrics: gRefCOCO [20], R-RefCOCO/+/g [30] and Ref-ZOM [11].
-Following previous GRES methods, we also conduct experiments on the RES dataset: RefCOCO/+/g [15, 27].
-Elaborate descriptions of each benchmark and metrics are in Sec.
-A.1, Sec.
-A.2.
-We adopt Swin Transformer [25] and BERT [5] as the backbone.
-Due to space limitations, complete implementation details are in Sec.
-A.3.
-Generalized Referring Expression Segmentation on gRefCOCO
+Finally, we convert the aggregated semantic-enriched segmentation $\mathcal{M}^{*}_{3}$ into a two-class mask $\mathcal{M}^{o}$ by multiplying kernel $B$ which is conditional on the $\mathcal{Q}_{3}^{L^{\prime}}$:
+$\mathcal{M}^{o}=\mathcal{M}^{*}_{3}\times B.$ (11)
 Method Backbone Validation Set Test Set A Test Set B
@@
 ReLA [22] Swin-T 56.87 57.73 44.07 65.29 66.02 48.68 57.37 56.86 50.31
-HDC (Ours) Swin-T 65.19 62.59 60.32 69.50 68.84 60.44 61.15 59.82 60.07
+CoHD (Ours) Swin-T 65.89 62.95 60.95 70.17 68.93 61.52 61.78 60.32 60.00
 LAVT [36] Swin-B 58.40 57.64 49.32 65.90 65.32 49.25 55.83 55.04 48.46
+DMMI [10] Swin-B 62.68 62.78 53.20 68.79 68.83 58.44 60.09 60.01 54.58
 ReLA [22] Swin-B 63.60 62.42 56.37 70.03 69.26 59.02 61.02 59.88 58.40
-HDC (Ours) Swin-B 68.28 65.42 63.38 72.52 71.60 65.29 63.85 62.79 60.68
+CoHD (Ours) Swin-B 68.42 65.17 63.68 72.67 71.85 64.00 63.60 62.63 60.37
 Table 1: Comparison with the state-of-the-art methods on gRefCOCO.
@@
 denotes that the results are from the official repository.
+* indicates that experimental results using the same hyperparameters configuration from the official repository.
 Method R-RefCOCO R-RefCOCO+ R-RefCOCOg
@@
 RefSegformer [33] 68.78 73.73 46.08 55.82 81.23 42.14 54.99 71.31 37.65
-HDC (Ours) 74.35 83.69 52.81 64.85 87.51 49.09 65.11 84.19 43.85
+CoHD (Ours) 74.16 84.27 53.61 64.59 87.49 49.07 63.56 82.68 42.16
 Table 2: Comparison with state-of-the-art methods on the R-RefCOCO/+/g dataset.
-4.2 Main Results
-Results on GRES tasks.
-To verify our effectiveness in the generalized setting, we first compare our method with previous specialist methods in Table 1 on gRefCOCO [20].
-It can be seen that our method achieves new state-of-art performance on all metrics of three evaluation sets in the large-scale GRES benchmark.
-Under fair comparison with the previous SOTA method ReLA [20], HDC achieves superior performance with a significant margin of $+4.7\%$, $+2.5\%$, $+2.8\%$ in gIoU of val, testA, testB set.
-Then, we report our results on Ref-ZOM benchmark [11](see in Table 3).
-As illustrated, our method outperforms all methods under a fair setting, e.g., $+6.3\%$ in Acc., $+1.1\%$ in oIoU.
-It is worth noting that our method is even better than GSVA [32], which utilizes Multi-Modal Large Language Model (MLLM) [23].
-In addition, we further make a comparison on R-RefCOCO/+/g [30].
-As depicted in Table 2, our method yields a significant improvement of $+6.7\%$, $+7.0\%$, $+6.2\%$ rIoU in R-RefCOCO/+/g compared with RefSegformer [30].
-Results on RES tasks.
-Following existing GRES methods [20, 11, 30], we train on each RES dataset without additional data [33].
-Due to the main focus on generalized formulation and space limitations, please see the Sec.
-B.1.
-HDC also outperforms all GRES methods in RES tasks.
-      Method       Backbone       Ref-ZOM Test Set
-      oIoU       mIoU       Acc.
-      MLLM Methods
-      LISA-V-7B [17]       SAM-ViT-H       61.46       60.14       72.58
-      GSVA-V-7B [32]       SAM-ViT-H       67.98       67.12       82.66
-      LISA-V-7B [17] (ft)       SAM-ViT-H       65.39       66.41       93.39
-      GSVA-V-7B [32] (ft)       SAM-ViT-H       68.13       68.29       94.59
-      Specialist Methods
-      MCN [26]       DarkNet-53       54.70       55.03       75.81
-      CMPC [12]       ResNet-101       55.72       56.19       77.01
-      VLT [6]       DarkNet-53       60.43       60.21       79.26
-      LAVT [33]       Swin-B       64.78       64.45       83.11
-      DMMI [11]       Swin-B       68.21       68.77       87.02
-      HDC (Ours)       Swin-B       69.31       68.81       93.34
-Table 3: Comparison with state-of-the-art methods on the Ref-ZOM dataset.
-HSD MAC gIoU cIoU N-acc.
-$53.52$ $56.55$ $41.11$
-$57.63$ $58.56$ $49.99$
-$63.15$ $60.67$ $55.84$
-$\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$
+3.3 Adaptive Object Counting
+Existing GRES methods tend to directly adopt a binary classification branch to judge the existence of objects.
+However, we observe that such a simple design exhibits less compatibility for general referring segmentation settings.
+Particularly, it roughly categorizes multiple-target and single-target scenarios into one objective, ignoring their inherent specificity, which can easily incur ambiguity when encountering implicit or counting expressions such as “a line of cars” or “two adult cow”.
+Meanwhile, several studies [23, 10] use the language reconstruction technique.
+Yet, it is not suitable in GRES since reconstructing unrelated content increases the burden of precise object awareness.
+Therefore, we propose Adaptive Object Counting (AOC), aiming to endow our model with category-specific counting ability in multiple/single/non-target situations.
+It significantly facilitates the precise understanding of both visual scenes and complicated sentences, i.e., counting, and compound structure expressions.
+Formally, the multi-modal semantic queries $\{\mathcal{Q}^{L^{\prime}}_{1},\mathcal{Q}^{L^{\prime}}_{2},\mathcal{Q}^{L^{\prime}}_{3}\}$ corresponding to each granularity are transformed into the counting vector of $C$ categories:
+$\mathcal{C}_{i}=\mathrm{MLP}(\mathcal{Q}^{L^{\prime}}_{i}),$ (12)
+where $\mathcal{C}_{i}\in\mathbb{R}^{N\times C}$.
+Then $\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\}$ are incorporated for the counting prediction $\mathcal{C}_{all}\in\mathbb{R}^{N\times C}$ by average pooling.
+Finally, we provide the explicit supervision on $\mathcal{C}_{all}$ by computing the $\mathit{L_{1}}$ distance between the multi-level accumulated results $\mathcal{C}^{pred}\in\mathbb{R}^{C}$ and the ground truth $\mathcal{C}^{gt}$, where $\mathcal{C}^{gt}$ (see in Fig.
+2) is annotated as the exact counts of each category.
+The whole process can be formulated as:
+[…]
+$\mathit{SmoothL1}(x)$ =$=\begin{cases}0.5\cdot x^{2}&\text{if }|x|<1.\\ |x|-0.5&\text{otherwise}.\end{cases}$
+On one hand, the AOC module is designed to seamlessly adapt to the specificity in generalized scenes, i.e., multiple/single/non-target tasks, which benefits the general object matching.
+On the other hand, guided by the the precise categories and counts of referent targets, our method facilitates comprehensive object perception in both image and language wise.
+3.4 Output and Loss
+[…]
+[…]
+Therefore, total loss with balance factor $\lambda$ is:
+[…]
+4 Experiments
+4.1 Experiment Setups
+We evaluate CoHD on three GRES benchmarks using official evaluation metrics: gRefCOCO [22], R-RefCOCO/+/g [33] and Ref-ZOM [10].
+Following previous GRES methods, we also conduct experiments on the RES dataset: RefCOCO/+/g [15, 29].
+Elaborate descriptions of each benchmark and metrics are in the supplementary material.
+We adopt Swin Transformer [27] and BERT [4] as the backbone.
+Due to space limitations, please defer to the supplementary material for complete implementation details.
+Methods Backbone RefCOCO RefCOCO+ RefCOCOg
+val test A test B val test A test B val${}_{\text{(U)}}$ test${}_{\text{(U)}}$
+Single Dataset (oIoU)
+MCN [28] Darknet53 62.44 64.20 59.71 50.62 54.99 44.69 49.22 49.40
+VLT [5] Darknet53 67.52 70.47 65.24 56.30 60.98 50.08 54.96 57.73
+ReSTR [16] ViT-B 67.22 69.30 64.45 55.78 60.44 48.27 - -
+CRIS [26] CLIP-R101 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
+LAVT [36] Swin-B 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
+RefSegformer [33] Swin-B 73.73 75.64 70.09 63.50 68.69 55.44 62.56 63.07
+RISCLIP [17] CLIP-B 73.57 76.46 69.76 65.53 70.61 55.49 64.10 65.09
+ReLA [22] Swin-B 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
+DMMI [10] Swin-B 74.13 77.13 70.16 63.98 69.73 57.03 63.46 64.19
+CoHD Swin-B 75.12 78.33 71.00 66.79 71.58 58.37 65.84 66.70
+Combined Dataset (mIoU)
+PolyFormer [25] Swin-B 75.96 77.09 73.22 70.65 74.51 64.64 69.36 69.88
+RISCLIP [17] CLIP-B 76.01 78.63 71.94 69.67 74.30 61.37 69.61 69.56
+CoHD Swin-B 78.11 80.39 75.20 72.03 76.37 65.45 70.83 72.11
+Table 3: Comparison with previous state-of-the-art GRES methods on classic RES benchmarks under different settings. indicates that the training includes additional datasets for pretraining.
+HSD AOC gIoU cIoU N-acc.
+53.52 56.55 41.11
+57.63 58.56 49.99
+63.15 60.67 55.84
+65.89 62.95 60.95
 Table 4: Effectiveness of the modules.
-STM STQ MMA gIoU cIoU N-acc.
-$57.63$ $58.56$ $49.99$
-$59.70$ $59.95$ $47.59$
-$62.06$ $62.36$ $52.30$
-$\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$
-Table 5: Effects of components of HSD module.
+Intra-Select Inter-Select gIoU cIoU N-acc.
+62.77 62.48 53.78
+63.56 62.32 56.16
+64.03 62.39 56.18
+65.89 62.95 60.95
+Table 5: Effects of components of DHA module.
+Method Backbone gIoU N-acc. Pr@0.7
+ReLA [22] Swin-T 70.30 44.07 52.21
+ReLA [22] Swin-B 73.47 56.37 59.89
+CoHD Swin-T 73.07 60.95 62.62
+CoHD Swin-B 75.78 63.68 66.24
+Table 6: Comprehensive evaluation.
+AOC Object Indicator gIoU cIoU N-acc. C-acc.
+binary head 63.15 60.67 55.84
+binary head 63.72 61.59 56.09 72.73
+count head 65.89 62.95 60.95 75.45
+Table 7: Impact of different strategies of object-existence identification.
+Mask Gen. Query Ref. DHA gIoU cIoU N-acc.
+63.31 62.74 54.15
+62.77 62.48 53.78
+65.89 62.95 60.95
+Table 8: Impact of different structure of HSD.
 Mask Decode gIoU cIoU
-Mask Attn. 63.36 61.97
 Deep Sup. 64.11 61.98
-MMA $\mathbf{65.19}$ $\mathbf{62.59}$
+DHA 65.89 62.95
 Both 63.35 62.34
 Table 9: Comparison of decoding methods.
-MAC Object Indicator gIoU cIoU N-acc. C-acc.
-binary head $63.15$ $60.67$ $55.84$
-binary head $63.72$ $61.59$ $56.09$ $72.73$
-count head $\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$ $\mathbf{74.58}$
-Table 7: Impact of different strategies of object-existence identification.
-Semantics Lang.Reactive gIoU cIoU N-acc.
-60.55 61.19 51.15
-63.31 61.56 56.10
-65.19 62.59 60.32
-Table 8: Impact of different structure of STQ.
+4.2 Main Results
+Figure 4: Segmentation Results. (a) and (b) are input images and segmentation results of CoHD with different referring expressions. The term Count specifies the output of AOC module.
+Figure 5: Visualization comparison in generalized setting. (a) and (b) are the segmentation results of ReLA and CoHD, respectively. The term Count specifies the output of AOC module.
+Results on GRES tasks.
+To verify the effectiveness of CoHD in the generalized setting, we first compare our method with previous specialist methods in Tab.
+1 on gRefCOCO [22].
+It can be seen that our method achieves new state-of-art performance on all metrics of three evaluation sets in the large-scale GRES benchmark.
+Under fair comparison with the previous SOTA method ReLA [22], CoHD achieves superior performance with a significant margin of $+4.8\%$, $+2.6\%$, $+2.6\%$ in gIoU of val, testA, testB set.
+Then, we report our results on R-RefCOCO/+/g [33].
+As depicted in Tab.
+2, our method yields a significant improvement of $+7.5\%$, $+6.9\%$, $+4.5\%$ rIoU in R-RefCOCO/+/g compared with RefSegformer [33].
+Due to the similarity of the datasets and space limitations, we report our results on Ref-ZOM benchmark [10] in the appendix.
+Our CoHD still outperforms all methods even GSVA [35], a MLLM based model, under a fair setting.
+Results on RES tasks.
+To fully illustrate the effectiveness of our design, We train CoHD with two settings [22, 10, 17] in Tab.
+3.
+CoHD exhibits significant superiority compared with previous methods.
+With the increment of the semantic context information brought by the data scale, the potential of HSD and AOC can be dramatically exploited.
 4.3 Ablation Study
-In this section, we conduct analysis experiments on the gRefCOCO [20] val set using Swin-Tiny as the visual backbone. Note that, we refer to additional ablation studies in Sec. B.2.
+In this section, we conduct analysis experiments on the gRefCOCO [22] val set using Swin-Tiny as the visual backbone. Note that, we refer to additional ablation studies in the appendix.
 Module effectiveness.
-We experiment with the effectiveness of the proposed Hierarchical Semantic Decoding (HSD) and Multi-Adaptive Object Counting (MAC) module.
-As shown in Table 8, With HSD hierarchically performing modality complementary and multi-level semantic aggregation, the performance is boosted about $4.1\%$ cIoU.
+We experiment with the effectiveness of the proposed Hierarchical Semantic Decoder (HSD) and Adaptive Object Counting (AOC) module.
+As shown in Tab.
+9, With HSD decouples the intricate referring semantics into the visual-linguistic hierarchy and dynamically aggregates with selective mechanism, the performance is boosted about $4.1\%$ cIoU.
 With the explicit counting ability, there is an improvement of $2\%$ cIoU.
@@
 Analysis of HSD.
-To investigate the validity of three key modules involved in HSD, we provide an analysis of the Semantic-to-Mask (STM), Semantic-to-Query (STQ), and Multi-level Mask Aggregation (MMA) components.
-As Table 8 shows, the mutual modality complementary for fine-grained semantic map generation at each level helps gain $2.1\%$ gIoU.
-Then, STQ formulates a higher level understanding of the linguistic information via the complementary of the referring holistic semantics, which brings a clear improvement of $2.3\%$ gIoU.
-since MMA aggregates across multi-level visual-linguistic semantics, the ablated of it leads to the degradation of $3.1\%$ gIoU.
+To investigate the validity of key modules involved in HSD, we provide an analysis of the Semantic Decoding Module (SDM) and Dynamic Hierarchical Aggregation (DHA) components.
+As Tab.
+9 shows, the Query Refinement which helps formulate a holistic understanding of the linguistic information with the assistance of the supervised semantic map, brings a clear improvement of $2.5\%$ gIoU.
+DHA is crucial in excavating the potential of hierarchical benefit, which gains $3.1\%$ gIoU.
 Variants of object-existence identification.
-We investigate the impact of different design variants on the object-existence justification that can be directly reflected by N-acc.
-As Table 8 shows, if maintaining the binary classification across all situations [20, 30], i.e., MAC is degraded to an auxiliary task, the counting ability still brings benefit.
-Furthermore, to explicitly quantify the counting ability, we calculate the accuracy between the prediction object count of each category and the true one, termed as C-acc.
-It can be seen that the simple binary head hampers the object perception, which results in $1.9\%$ C-acc drop.
-With the subsequent design of counting head upon MAC which facilitates the comprehensive object perception under each situation, HDC yields the most robust performance.
-Analysis of STQ Structure.
-Results on Table 8 demonstrate that the incorporated referring semantics at the query level is crucial, which contributes to $+2.8\%$ gIoU.
-Moreover, since the subsequent language reactivation prevents the shifted emphasis of key linguistic contents and formulates a deeper understanding of them via the semantics supplement, the absence of it causes $1.9\%$ gIoU degradation.
+We investigate the impact of different variants on the object-existence justification that N-acc can directly reflect.
+As Tab.
+9 shows, if maintaining the binary classification across all situations [22, 33], i.e., AOC degraded to an auxiliary task, the counting ability still benefits.
+We adopt the accuracy between the prediction object count of each category and the true one, termed as C-acc to quantify the counting ability.
+It can be seen that the simple binary head hampers the object perception, which results in $2.7\%$ C-acc drop.
+With the subsequent design of counting head upon AOC which facilitates the comprehensive object perception under each situation, CoHD yields the most robust performance.
+Analysis of intra- and inter-selection.
+Results on Tab.
+9 demonstrate that both intra-selection and inter-selection mechanisms are indispensable components in DHA, since the former helps enhance the related region activation at the token level and the latter emphasizes the saliency at the granularity level.
+The ablation of each will lead to $1.8\%$ and $2.3\%$ degradation respectively.
 Mask decoding method.
-To prove the effectiveness of MMA, we compare different prevalent mask decoding methods.
-As shown in Table 8, replacing MMA with mask attention [2] mechanism at each level to generate the final mask prediction is inferior.
-This can be attributed to the restriction on specific regions at the preliminary level is susceptible to the following modality interaction.
-Moreover, we observe that the utilization of deep supervision at each level with or without aggregation falls below anticipated outcomes.
-We speculate that it deteriorates the nature of the hierarchical structure that target responses from levels are different but adaptively reciprocal.
+To prove the effectiveness of DHA, we compare different prevalent mask decoding methods.
+As shown in Tab.
+9, we observe that the utilization of deep supervision at each level with or without aggregation falls below anticipated outcomes.
+We speculate that it deteriorates the hierarchical nature that object responses from levels are different but adaptively reciprocal.
 Comprehensive evaluation in generalized settings.
-Figure 4: Comprehension evaluation from three perspectives.
-To quantify our overall superiority in addressing the challenges introduced in GRES, we synthetically compare HDC with the previous SOTA method, ReLA [20] from three perspectives.
-The horizontal axis of Fig.
-4 denotes the gIoU obtained from only Multiple-target scenarios, the vertical axis indicates the N-acc.
-calculated in Zero-target cases individually, while the number inside the circle represents the total Pr@0.7 of the One-target situation.
-As illustrated, HDC shows absolute dominance across dimensions compared to ReLA.
-Figure 5: Segmentation Results. (a) and (b) are input images and segmentation results of our HDC with different referring expressions. The term Count specifies the output of MAC module.
-Figure 6: Visualization comparison in generalized setting. (a) and (b) are the segmentation results of ReLA and our HDC, respectively. The term Count specifies the output of MAC module.
+To quantify our overall superiority in addressing the challenges introduced in GRES, we synthetically compare CoHD with the previous SOTA method, ReLA [22] from three perspectives: 1) gIoU, obtained from only multiple-target scenarios.
+2) N-acc., calculated in non-target cases individually.
+3) Pr@0.7, measured by the single-target situation.
+As illustrated in Tab.
+9, CoHD shows absolute dominance across dimensions compared to ReLA.
 4.4 Qualitative Results
-Benefiting from the hierarchical semantic aggregation with well-aligned granularities and the counting ability, Fig.
+Benefiting from the dynamic hierarchical aggregation on well-aligned semantics of different granularity and the counting ability, Fig.
 4 and Fig.
 5 demonstrate our superiority in processing the complex spatial relationship between instances and intricate referring text in multiple/single/non-target scenarios.
-In contrast, ReLA [20] fails to adequately address these challenges in GRES due to its incomplete modeling of visual-linguistic correlation.
-More visualization results are in Appendix C.
+In contrast, ReLA [22] fails to adequately address these challenges in GRES due to its incomplete multi-granularity object information modeling.
+More visualization results are in the supplementary material.
 5 Conclusion
-In this paper, we propose a framework called HDC for GRES to boost the multi-grained comprehension of complex referring semantic associations at different levels.
-With progressively transferring modality complementary information across granularities, HDC aggregates each visual-linguistic correspondence for multi-level semantic decoding.
-Moreover, with well-aligned semantic contexts, we empower HDC with category-specific counting ability to promote comprehensive object perception in multiple/single/non-target scenarios.
-Experimental results show the dominance of HDC in all benchmarks and the superiority in addressing the new challenges of GRES.
+This paper proposes a framework called CoHD for GRES, aiming to decouple the complex referring semantics into different levels for boosting multi-granularity comprehension.
+With the well-aligned visual-linguistic hierarchy of different granularity, CoHD dynamically aggregates it with the intra- and inter-selection.
+Moreover, with complete semantic context, we empower CoHD with category-specific counting ability to promote the overall object perception in multiple/single/non-target scenarios.
+Experimental results on GRES benchmarks show the effectiveness of CoHD and the superiority in addressing the new challenges.
 Appendix
-Appendix A Additional Details on Experiment Setup
-A.1 Datasets
+6 Additional Details on Experiment Setup
+6.1 Datasets
 gRefCOCO.
@@
 RefCOCO [15], RefCOCO+ [15], and RefCOCOg [29] are three standard RES benchmarks, each of which contains 19,994, 19,992, and 26,711 images, with 50,000, 49,856, and 54,822 annotated objects and 142,209, 141,564, and 104,560 annotated expressions, respectively.
-A.2 Metrics
+6.2 Metrics
 For GRES, following [22], we measure the effectiveness of our model by Pr@0.7, gIoU, cIoU and N-acc for gRefCOCO.
@@
 Similarly, mRR in R-RefCOCO calculates the empty-target expression recognition rate of each image and averages these across the entire dataset.
-A.3 Implementation Details
+6.3 Implementation Details
+Experiment setup.
 […]
@@
 All experiments are conducted with 8×8$\times$8 × A10 and each takes up about one day with 13GB $\sim$ 18GB memory occupied i.e., it depends on the backbone.
-Appendix B Additional Experiments
-B.1 Performance on Classic RES Task
-As shown in Table 9, our performance is higher than the previous state-of-art GRES methods with a consistent improvement of $1\sim 3\%$ on three datasets, showing that the multi-level semantic aggregation with modality alignment at each granularity and the category-specific object counting capability remains effective in the simpler setting.
-Methods Backbone RefCOCO RefCOCO+ RefCOCOg
-val test A test B val test A test B val${}_{\text{(U)}}$ test${}_{\text{(U)}}$
-MCN [26] Darknet53 62.44 64.20 59.71 50.62 54.99 44.69 49.22 49.40
-VLT [6] Darknet53 67.52 70.47 65.24 56.30 60.98 50.08 54.96 57.73
-ReSTR [16] ViT-B 67.22 69.30 64.45 55.78 60.44 48.27 - -
-CRIS [24] CLIP-R101 70.47 73.18 66.10 62.27 68.08 53.68 59.87 60.36
-LAVT [33] Swin-B 72.73 75.82 68.79 62.14 68.38 55.10 61.24 62.09
-RefSegformer [30] Swin-B 73.73 75.64 70.09 63.50 68.69 55.44 62.56 63.07
-ReLA [20] Swin-B 73.82 76.48 70.18 66.04 71.02 57.65 65.00 65.97
-DMMI [11] Swin-B 74.13 77.13 70.16 63.98 69.73 57.03 63.46 64.19
-Ours Swin-B 74.81 78.32 70.95 66.80 71.88 58.90 65.59 66.40
-Table 9: Comparison with previous state-of-the-art GRES methods on classic RES benchmarks.
+Counting labels formulation.
+We elaborate on the formulation of the counting label.
+All mentioned datasets adhere to the COCO  [20] annotation format.
+On the one hand, for RES datasets, each annotated expression is accompanied by a label (category_id) corresponding to the target category.
+On the other hand, a list of target categories is incorporated with the additional multi-target scenario for GRES dataset.
+Consequently, the count of objects can be derived from the number of categories and each classification label is from the given category of the annotated object.
+It signifies that the construction of counting ground truth $\mathcal{C}^{gt}$ is straightforward and the additional information extraction is unnecessary.
+Note that, taking the long-tail distribution of the 80 original COCO categories into consideration where most of the categories are annotated as 00, we instead utilize the 12 super-categories to narrow down the referential search space for providing more precise supervision.
+Method Backbone Ref-ZOM Test Set
+mIoU oIoU Acc.
+MLLM Methods
+LISA-V-7B [18] SAM-ViT-H 61.46 60.14 72.58
+GSVA-V-7B [35] SAM-ViT-H 67.98 67.12 82.66
+LISA-V-7B [18] (ft) SAM-ViT-H 65.39 66.41 93.39
+GSVA-V-7B [35] (ft) SAM-ViT-H 68.13 68.29 94.59
+Specialist Methods
+MCN [28] DarkNet-53 54.70 55.03 75.81
+CMPC [11] ResNet-101 55.72 56.19 77.01
+VLT [5] DarkNet-53 60.43 60.21 79.26
+LAVT [36] Swin-B 64.78 64.45 83.11
+DMMI [10] Swin-B 68.21 68.77 87.02
+CoHD (Ours) Swin-B 69.81 68.99 93.34
+Table 10: Comparison with state-of-the-art methods on the Ref-ZOM dataset.
 $\lambda_{mask}$ $\lambda_{count}$ gIoU cIoU N-acc. C-acc.
@@
 $5$ $0.1$ 64.15 62.54 57.24 71.46
-$2$ $0.1$ $\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$ $\mathbf{74.58}$
+$2$ $0.1$ 65.89 62.95 60.95 75.45
 Table 11: Results of different ratios of loss.
-$\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ gIoU cIoU N-acc. C-acc.
-$0.1$ $1.0$ $1.0$ $64.34$ $62.27$ $58.57$ $72.98$
-$1.0$ $0.2$ $1.0$ $63.62$ $61.96$ $55.99$ $73.50$
-$1.0$ $1.0$ $0.2$ $63.11$ $62.25$ $54.32$ $72.95$
-$0.2$ $0.2$ $1.0$ $62.79$ $62.28$ $54.04$ $70.93$
-$1.0$ $1.0$ $1.0$ $\mathbf{65.19}$ $\mathbf{62.59}$ $\mathbf{60.32}$ $\mathbf{74.58}$
-Table 11: Results of different aggregation degrees among levels in MMA.
-B.2 Additional Ablation Studies
+7 Additional Experiments
+7.1 Discussion on Inconsistent Performance
+As shown in the manuscript, we notice that the performance improvement is inconsistent between RES and GRES.
+We believe it can be attributed to two folds: 1) Restricted effectiveness on the generalized design: It is obvious that the diversity of referring scenarios in GRES is much more plentiful compared to the RES since RES only includes one-to-one referent case.
+Due to the inadequate referring semantics between visual and linguistic and the lack of enriched contextual information, e.g., Spatial relationship between instances, counting, or compound structure expressions, it is believed that the great potential of HSD is constrained.
+Moreover, the effectiveness of object counting mechanism in RES is also underestimated.
+In GRES, it seamlessly integrates all specificities of each scenario by embodying each case into count- and category-level supervision.
+In terms of RES, the simplification of the formulation limits the potential of object counting, where count number supervision is missing.
+2) Imbalance dataset scale: The samples of GRES dataset gRefCOCO is 230,944 samples, while that of each RES dataset is: RefCOCO: 120,624 RefCOCO+: 120,191 RefCOCOg: 80,544.
+That indicates the specific design and effort on hyper-parameter studies in RES weighs more crucial.
+Since our main focus is on the generalized referring which accompanies more useful real-world applications, complex hyperparameter studies on RES are not applied.
+Considering the hypothesis mentioned above, we believe that enlarging the scale of the dataset can partly alleviate the phenomenon.
+The results in the joint dataset training [17] demonstrate that our CoHD brings more improvements compared with the single scenario, showing that the potential of our generalized paradigms can be exploited with enriched contextual information.
+7.2 Performance on Ref-ZOM Dataset
+We report our results on Ref-ZOM benchmark [10] in Tab.
+10.
+As illustrated, our method outperforms all methods under a fair setting, e.g., $+6.3\%$ in Acc., $+1.6\%$ in mIoU.
+It is worth noting that our method is better than GSVA [35], which utilizes Multi-Modal Large Model (MLLM) [24].
+Method Backbone Parameters T-FLOPs gIoU cIoU N-acc.
+ReLA Swin-T 163M 0.066T 56.87 57.73 44.07
+CoHD Swin-T 185M 0.068T 65.89 62.95 60.95
+DMMI Swin-B 341M 0.392T 62.68 62.77 53.20
+ReLA Swin-B 226M 0.131T 63.60 62.42 56.37
+CoHD Swin-B 248M 0.133T 68.42 65.17 63.68
+Table 12: Performance and efficiency comparison with previous SOTA method under different backbones.
+Stacked Layer gIoU cIoU N-acc.
+2 66.33 64.58 58.48
+3 68.42 65.17 63.68
+4 66.70 64.94 60.49
+Table 13: Impact on different numbers of stacked layers in the Semantic Decoding Module.
+7.3 Additional Ablation Studies
 Loss ratio.
 […]
-Different Aggregation Degrees.
-The $\alpha_{i}$ controls the fusion ratio between levels in Multi-level Mask Aggregation module, which represents the importance of each level in the final mask generation.
-Therefore, We investigate the effect of it in Table 11.
-As illustrated, ignorance of any level will lead to inferior performance.
-We speculate that the complex spatial relationship between instances with the intricate referring expression highlights the benefits of hierarchical structures where information in each level is adaptively reciprocal.
-Appendix C Visualizations.
-Figure 7: Visualization in Multi-Level Mask Aggregation.
-(a) and (c) indicate the input image and corresponding segmentation result of our HDC, respectively.
-(b) illustrates the aggregated activation map at each level.
-C.1 Multi-level Mask Aggregation Visualization.
-As illustrated in the main paper, we hierarchically aggregate each visual-linguistic correspondence in different granularity for multi-level semantic decoding.
+Stacked layer in Semantic Decoding Module.
+We experiment with the impact on the number of stacked layers in the semantic decoding module.
+As shown in Tab.
+13, the insufficient or excessive layers both lead to decreased performance due to incomplete semantic context modeling or over-exaggeration of the redundant contents.
+Efficiency v.s. performance.
+We provide detailed comparisons (including model parameters and T-FLOPs.) on the gRefCOCO val set and the results are shown in Tab.
+12.
+It can be seen that our CoHD-B outperforms previous SOTA methods ReLA-B by 4.8% gIoU and 2.7% cIoU with slight parameters increased, i.e., 22M, and even no T-FLOPs costs.
+It is worth noting that the performance of CoHD-T is even better than ReLA-B with fewer parameters.
+In addition, the number of parameters consumed by HSD is 14.7% of the whole model, but with significant performance improvement.
+8 Visualizations.
+8.1 Multi-granularity Mask Aggregation Visualization
+As illustrated in the main paper, we hierarchically aggregate each visual-linguistic correspondence in different granularity for hierarchical semantic decoding.
 To further prove the rationality of our design, we visualize the aggregated activation map (originally from the semantic map) at each level in Fig.
-7.
+6.
 As observed, with progressive mutual modality complementary across granularities, the desired regions of the referent target are fully activated with the increase of the levels.
-C.2 Additional Segmentation Comparisons.
-As illustrated in the main paper and Fig.
-6, HDC can better handle the extreme challenges of GRES under multiple/single/empty target scenarios compared with the previous SOTA method ReLA [20].
+8.2 Additional Segmentation Comparisons
+As illustrated in the main paper, CoHD can better handle the extreme challenges of GRES under multiple/single/empty target scenarios compared with the previous SOTA method ReLA [22].
 Here, we incorporate more cases in Fig.
 8 to demonstrate our superiority in modeling the intricate referring association between visual and linguistic.
-C.3 Segmentation Results of HDC.
+8.3 Segmentation Results of CoHD
 Fig.
-8 demonstrates the effectiveness of HDC in complex generalized settings, e.g., complex geometrical relationships between instances, deceptive objects, compound sentence structures, and intricate associations between referring expressions and images.
-Appendix D Limitations.
-Benefiting from the hierarchical semantic decoding and the explicit counting ability, our method HDC sufficiently addresses the limitations of existing GRES and achieves superiority in meeting the challenges of GRES.
+7 demonstrates the effectiveness of CoHD in complex generalized settings, e.g., complex geometrical relationships between instances, deceptive objects, compound sentence structures, and intricate associations between referring expressions and images.
+9 Limitations
+Benefiting from the more compatible design, i.e., hierarchical semantic decoding paradigm and the explicit counting ability, our method CoHD sufficiently addresses the limitations of existing GRES and achieves superiority in meeting the challenges of GRES.
 However, there are some potential limitations.
@@
 Although we compensate it with sentence-level textual features, it still remains unsatisfactory in terms of the incompleteness of fine-grained target description.
-We think that how to fully utilize the textual expression in GRES is an interesting future direction.
-Appendix E Error Bar.
-We have retrained our model several times on gRefCOCO [20]. The results demonstrate that the randomness of the model has little effect on the performance, i.e., the standard variation is about $0.36$.
-Appendix F Broader Impacts.
-Malicious use of the GRES model may lead to potential negative societal impacts, including but not limited to unauthorized surveillance or privacy-infringing tracking.
-However, we firmly believe that the task itself is neutral with positive implications, such as image editing and human-robot interaction.
-Figure 8: Segmentation results of our method in generalized settings.
-(a) denotes the input image, and (b) showcases the segmentation results of HDC under multiple/single/non-target situations with different referring expressions.
+We believe that how to fully utilize the textual expression in GRES is an interesting future direction.
+Figure 6: Visualization in Multi-granularity Mask Aggregation.
+(a) and (c) indicate the input image and corresponding segmentation result of our CoHD, respectively.
+(b) illustrates the aggregated activation map at each granularity.
+Figure 7: Segmentation results of CoHD in generalized settings.
+(a) denotes the input image, and (b) showcases the segmentation results of CoHD under multiple/single/non-target situations with different referring expressions.
 Figure 8: Segmentation results comparison in the generalized setting.
-(a) denotes the input image, (b) and (c) are segmentation results of ReLA and our HDC, respectively.
+(a) denotes the input image, (b) and (c) are segmentation results of ReLA and CoHD, respectively.
 Empty, Single, and Multiple target referent situations with different referring expressions are in top-to-bottom accordance.